### v2

# Maximum Implementable Improvements over Deferred Acceptance

Abstract
The Deferred Acceptance (DA) mechanism produces stable but Pareto-inefficient school assignments. While mechanisms like Efficiency-Adjusted DA address this inefficiency by asking students to waive priorities at unattainable schools, they can leave significant gains unrealized—sometimes improving only two students when all but one could benefit. Surprisingly, such mechanisms can be doubly dominated, creating fewer improvements while generating more justified envy.
We introduce the Maximum Implementable Improvement over the DA mechanism (MIIDA), which maximizes the number of students made better off given priority-waiving consents. MIIDA weakly Pareto dominates DA, is never doubly dominated, and incentivizes students to consent.
school choice,
maximum improvements,
consent,
C78,
D47,
keywords:
keywords:
[class=JEL]
[…]
Knipe: tknipe01@qub.ac.uk, Ortega: j.ortega@qub.ac.uk. We acknowledge helpful discussions with Pablo Arribillaga, Péter Biró, Battal Doğan, Martin Meier, and Gabriel Ziegler. T. Knipe acknowledges financial support for doctoral studies from Northern Ireland’s Department for the Economy.
[…]
1 Introduction
The student-proposing Deferred Acceptance (DA) algorithm has become a pillar of market design, combining elegant theoretical properties with widespread adoption in school assignment systems worldwide. While DA guarantees stable matchings, it can produce allocations that are Pareto-inefficient for students in theory and in practice. This inefficiency has motivated the development of mechanisms that weakly Pareto-dominate DA by implementing mutually beneficial trades among students, which only violate the priorities of students who willingly consent to such violations. The most prominent solutions—Efficiency-Adjusted DA (EADA, Kesten 2010) and Top Trading Cycles applied to DA assignments (DA+TTC)—fix DA’s inefficiency without harming any student while, at the same time, preserving weak but meaningful stability and incentive properties. In particular, EADA has shown a remarkable performance in both theory and the lab, and has been considered for practical implementation in Flanders (discussed in Section 2).
In this paper, we analyze how many students can improve their DA assignment through the trading cycles implemented by EADA and DA+TTC, while tracking the number of blocking pairs these exchanges introduce.
We find that both mechanisms can perform poorly when judged solely by the number of students they improve. In problems with $n$ students, we can construct examples where $n-1$ students could benefit. However, EADA and DA+TTC forgo this large cycle involving almost all students and perform a single two-way exchange instead. Therefore, the improvement ratio of both mechanisms, which tells us how many efficiency swaps they miss, is $\frac{n-1}{2}$ (the largest possible ratio obtained by an efficient mechanism that improves upon DA) and grows linearly with the size of the school choice problem (Theorem 1).
Surprisingly, we find that existing mechanisms can be doubly dominated, simultaneously improving fewer students while generating more blocking pairs in specific problems than alternative allocations (Proposition 1). This counterintuitive result reveals an unexploited margin on the efficiency-stability frontier: some Pareto improvements strictly dominate EADA and DA+TTC on both dimensions, yet mechanisms that achieve this have not been explored in the literature. These larger improvements can sometimes require fewer consents overall, and specifically fewer non-beneficiary consents—agreements from students who do not directly benefit from the proposed trades and must therefore act altruistically—compared to the smaller improvements implemented by existing mechanisms (Proposition 2).
Furthermore, EADA and DA+TTC prove incomparable regarding double dominance—in different problem instances, each can produce matchings that doubly dominate the other. This mutual non-dominance suggests that the choice between EADA and DA+TTC involves case-specific tradeoffs rather than universal superiority.
The potentially small number of improvements carried out by existing mechanisms motivates us to propose a new mechanism that always improves the placement of as many students as possible. This algorithm computes DA and, based on the corresponding envy digraph, chooses the collection of disjoint trading cycles that maximizes the number of students involved given the consent structure available. We call this mechanism Maximum Implementable Improvement over Deferred Acceptance (MIIDA). MIIDA weakly Pareto dominates DA and is Pareto-efficient when all students consent. It can generate less or more justified envy, depending on the problem, but it is never doubly dominated because it improves as many students as possible. With regard to incentives, it gives every student a weak incentive to consent to waive their priorities. Finally, it can be computed in polynomial time (Theorem 2).
While MIIDA provides valuable theoretical insights about mechanism performance, it has important drawbacks too: it can be manipulated and can generate more blocking pairs than EADA and DA+TTC. Moreover, the set of blocking pairs that it generates may be a strict superset of the blocking pairs generated by other setwise justified envy minimal mechanisms such as EADA. Such negative properties are unavoidable: any efficient mechanism that weakly Pareto dominates DA and is setwise justified envy minimal has an improvement ratio $\frac{n-1}{2}$, leaving large efficiency gains unrealized (Theorem 3).
Our findings show that consent structures already embedded in existing mechanisms contain unexploited potential for achieving greater welfare improvements, as larger improvements can sometimes be achieved while requiring fewer consents overall, and particularly fewer altruistic consents from non-beneficiary students. Moreover, these larger improvements can occasionally generate fewer blocking pairs than smaller ones, creating a counterintuitive bidimensional improvement on the efficiency-stability trade-off. Overall, MIIDA serves as a useful benchmark that allows us to quantify precisely how much improvement potential is foregone by existing mechanisms and reveals opportunities for further mechanism design in this space.
To summarize, our findings contribute to the literature in three ways. First, we provide tight bounds on the improvement ratios of existing mechanisms, showing that both EADA and DA+TTC can leave significant efficiency gains unrealized in worst-case scenarios, and show that these efficiency gains may be compatible with parallel stability gains. Second, we introduce MIIDA as a new mechanism in the relatively under-explored class of efficient mechanisms that Pareto-dominate DA and document its properties and weaknesses. Third, we establish an impossibility result: any mechanism that attains a non negligible improvement size cannot be justified envy minimal.
Outline.
The remainder of this paper is organised as follows. Section 2 discusses the literature. Section 3 introduces the model. Section 4 presents the limitations of existing mechanisms. Section 5 introduces the Maximum Implementable Improvement over Deferred Acceptance mechanism. Section 6 concludes.
2 Related Literature
DA’s Inefficiency.
The Pareto-inefficiency of the student-proposing Deferred Acceptance algorithm has been recognized since the early studies on matching markets and school choice (Roth, 1982; Abdulkadiroğlu and Sönmez, 2003). Kesten (2010) provided the first theoretical bounds on DA’s inefficiency, establishing that the welfare losses can be substantial.
DA’s inefficiency has been consistently documented in practice. Studies including Abdulkadiroglu et al. (2005), Abdulkadiroğlu et al. (2009), Che and Tercieux (2019), Abdulkadiroğlu et al. (2020), and Ortega and Klein (2023) consistently find that DA leaves significant welfare gains on the table across different school choice systems. On the other hand, theoretical literature has identified restrictive conditions on preference and priority structures under which DA achieves efficiency (e.g. Ergin, 2002; Cantillon et al., 2024).
Mechanisms That Improve Upon DA.
The most prominent solution to DA’s inefficiency is the Efficiency-Adjusted Deferred Acceptance mechanism, proposed in a seminal paper by Kesten (2010). EADA identifies Pareto-improving trading cycles among students that only violate the priorities of students who benefit from consenting to waive their priorities, ensuring such students have no incentive to withhold consent. While Kesten notes that EADA’s improvements may not be the largest possible and argues that selecting larger cycles could be problematic due to the greater number of consents required, our contribution is to quantify exactly how many improvements EADA may be missing and to show that such larger improvements need not always require more consent nor generate more justified envy.
EADA has been extensively studied both theoretically and experimentally, with the literature consistently documenting its strong performance across multiple stability and incentives criteria (Bando, 2014; Tang and Yu, 2014; Dur et al., 2019; Troyan et al., 2020; Troyan and Morrill, 2020; Ehlers and Morrill, 2020; Tang and Zhang, 2021; Reny, 2022; Chen and Möller, 2023; Cerrone et al., 2024; Cheng et al., 2024; Shirakawa, 2025).
Form this literature, the papers closer to ours are Doğan and Ehlers (2021a) and Kwon and Shorrer (2020). These papers show that EADA (among other mechanisms) is justified envy minimal among efficient mechanisms in a set inclusion sense using blocking pairs or triplets, respectively. [^1] Abdulkadiroğlu et al. (2020) and Doğan and Ehlers (2021b) show that the Top Trading Cycles algorithm is also (setwise) justified envy minimal. While this property has made EADA attractive for practical implementation, we prove that any mechanism satisfying setwise justified envy minimality must have an improvement ratio of at least $\frac{n-1}{2}$, leaving large welfare gains unrealized in specific problems. Our MIIDA mechanism deliberately relaxes this constraint to maximize the number of students who benefit, demonstrating that the choice between minimal justified envy and maximal improvement size represents a stark trade-off.
Another natural approach to improving DA is to apply the Top Trading Cycles mechanism to DA’s allocation. While this mechanism is know in the literature, it has received surprisingly limited formal analysis (with some exceptions, see Alcalde and Romero-Medina 2017; Kesten 2010; Troyan et al. 2020).
The broader literature has explored various other approaches to improving DA. Kesten and Kurino (2019) examine conditions under which DA can be improved while maintaining strategy-proofness, while Doğan and Yenmez (2020) study consistency properties of DA improvements. Dur and Morrill (2020) examine the equilibrium performance of mechanisms that dominate DA. Ortega et al. (2025) demonstrate that certain limitations of DA—including rank-inefficiency, inequality and segregation—cannot be addressed by any Pareto-improvement.
Scope of Efficiency Gains.
Ortega et al. (2024) show that DA is Pareto-inefficient with high probability in large i.i.d. random markets, and that the envy digraph typically contains a giant strongly connected component, implying that most students can potentially benefit from disjoint trading cycles. Furthermore, all efficient mechanisms that dominate DA tend to improve nearly every student. This finding provides context for our analysis. While their paper shows that improvement mechanisms converge in their scope in large i.i.d. random markets, our paper demonstrates that this convergence does not hold in specific instances. We show that significant gaps can emerge between mechanisms in terms of both the number of students improved and the stability costs incurred.
3 Model
Following Abdulkadiroğlu and Sönmez (2003), a school choice problem $P$ consists of a set of $n$ students $I$ and a set of schools $S$. Each student $i\in I$ has a strict preference relation $\succ_{i}$ over the schools. Each school $s\in S$ has a quota of available seats $q_{s}$ and a strict priority relation $\triangleright_{s}$ over the students. The school $s_{\emptyset}$ with $q_{s_{\emptyset}}=\infty$ denotes being unassigned. For a given school choice problem $P$, a matching $\mu$ is a mapping from $I$ to $S$ such that no school is matched to more students than its quota. We denote by $\mu_{i}$ the school to which student $i$ is assigned.
The function $\textup{rk}_{i}:S\rightarrow\{1,\ldots,n\}$ specifies the rank of school $s$ according to the preference relation $\succ_{i}$ of student $i$:
$\textup{rk}_{i}(s)=|\{s^{\prime}\in S:s^{\prime}\succ_{i}s\}|+1,$ (1)
so that student $i$’s most preferred school gets a rank of $1$. With some abuse of notation, we use the same rank function to specify the students’ rank per the priority profile of schools.
Properties.
A matching $\mu$ weakly Pareto-dominates matching $\nu$ if, for every student $i\in I$, $\textup{rk}_{i}(\mu_{i})\leq\textup{rk}_{i}(\nu_{i})$. A matching $\mu$ Pareto-dominates matching $\nu$ if $\mu$ weakly dominates $\nu$ and there exists a student $j\in I$ with $\textup{rk}_{j}(\mu_{j})<\textup{rk}_{j}(\nu_{j})$. A matching is Pareto-dominated if there exists a matching that Pareto-dominates it and is Pareto-efficient if it is not Pareto-dominated.
Student $i$ desires school $s$ in matching $\mu$ if $\textup{rk}_{i}(s)<\textup{rk}_{i}(\mu_{i})$ and he envies student $j$ at matching $\mu$ if $\textup{rk}_{i}(\mu_{j})<\textup{rk}_{i}(\mu_{i})$. We say that student $j$ violates student $i$’s priority at school $s$ in matching $\mu$ if $i$ desires $s$, $\mu_{j}=s$, and $\textup{rk}_{s}(i)<\textup{rk}_{s}(j)$. If so, we call ($i,j,s$) and ($i,s$) a blocking triplet and a blocking pair at matching $\mu$, respectively. A matching $\mu$ is non-wasteful if every school $s$ that is desired by some student in $\mu$ satisfies $|\mu^{-1}_{s}|=q_{s}$. A matching $\mu$ is stable if it is non-wasteful and no student’s priority at any school is violated in $\mu$.
Mechanisms.
A mechanism is a function that maps every school choice problem to a matching. We particularly focus on three mechanisms.
  1. 1.
    Deferred Acceptance (DA). The student-proposing DA mechanism (Gale and Shapley, 1962) works as follows. Students apply to their most preferred school. Schools tentatively accept applicants up to their capacity, rejecting the rest. Rejected students apply to their next choice in subsequent rounds. The algorithm terminates when no new rejections occur, yielding the student-optimal stable matching $\mathtt{DA}(P)$.
  2. 2.
    Efficiency-Adjusted DA (EADA). To execute EADA, first we define an interrupter, which is a student-school pair $(i,s)$ such that:
    • student $i$ is tentatively accepted at $s$ in step $t$,
    • student $j\neq i$ is rejected at $s$ in step $t^{\prime}\geq t$,
    • student $i$ is rejected from school $s$ in step $t^{\prime\prime}\geq t^{\prime}$
    To compute EADA, we follow the following procedure:
    1. (a)
      Run DA, find the last interrupter ($i,s$), in which student $i$ consents to waive his priority at school $s$.
    2. (b)
      Eliminate $s$ from the list of preferences of student $i$ and re-run DA.
    3. (c)
      Find the last consenting interrupter. If any, go to step 2. If there are no interrupters, the last DA allocation is implemented.
    When all interrupter students consent to waive their priorities, EADA is Pareto-efficient. In that case, EADA can alternatively be computed by running DA, finding schools that did not reject anyone, matching them permanently to their DA match and removing such students and schools, and re-executing DA on the smaller problem until all schools are rejected (Tang and Yu, 2014).
  3. 3.
    Top Trading Cycles on top of DA (DA+TTC) This mechanism first computes DA, then applies Gale’s Top Trading Cycles algorithm where each student initially "owns" their DA assignment (Shapley and Scarf, 1974). Students point to owners of their most preferred schools, and all students in cycles trade. The process repeats until no cycles remain.
Envy Digraphs and Improvement Cycles.
To analyse improvements over DA, we use the envy digraph proposed by Ortega et al. (2024), which illustrates who envies whom after the initial DA allocation (envy need not be justified).
Definition 1.
The envy digraph $G^{\mathtt{DA}(P)}$ is a directed graph where:
  • Nodes represent students in $I$.
  • A directed edge $(i,j)$ exists if student $i$ envies student $j$’s assignment under DA, i.e. if $\textup{rk}_{i}(\mathtt{DA}_{j}(P))<\textup{rk}_{i}(\mathtt{DA}_{i}(P))$.
We denote the edge $(i,j)$ as $i\rightarrow j$. A cycle in the digraph is a sequence of nodes $(i_{0}\to i_{1}\to\ldots\to i_{k}\to i_{0})$ such that there is a directed edge between each pair of consecutive nodes and no edge is repeated. A trading cycle is a cycle in which every node appears exactly once, except for $i_{0}$. Unless specifically stated otherwise, every cycle mentioned henceforth is a trading cycle.
As argued in Section 2, DA is rarely Pareto-efficient. Therefore, we focus on school choice problems where $G^{\mathtt{DA}(P)}$ contains at least one cycle. To systematically identify and implement improvements over DA, we use the concept of a feedback set.
Definition 2.
A feedback set $F$ is a collection of pairwise-disjoint cycles in $G^{\mathtt{DA}(P)}$ such that, when all nodes in these cycles are removed (together with their adjacent edges), the resulting subgraph is acyclic. [^2] Feedback node sets are sometimes called feedback vertex sets (Karp, 2010). The concept of a “cycle cover” is related, but requires that every node must be part of a cycle. The problem of choosing the maximum feedback set is sometimes known as cycle packing.
Feedback sets represent sets of disjoint trading cycles that can be simultaneously implemented to improve upon DA. We use $V(F)$ to denote the set of nodes (students) that are included in some cycle in $F$. If $i\in V(F)$, we say that $F$ covers $i$. A feedback set $F^{\star}$ is maximum if, for any other feedback set $F^{\prime}$, we have $|V(F^{\prime})|\leq|V(F^{\star})|$.
Given a feedback set $F$, we define the matching $\mu^{F}$ as the allocation obtained when the trades in each cycle of $F$ are implemented, starting from DA. Formally:
  • For student $i\in V(F)$ belonging to cycle $C$, $\mu^{F}_{i}=\mathtt{DA}_{\text{succ}(i)}(P)$, where $\text{succ}(i)$ denotes the successor of $i$ in cycle $C\in F$.
  • For student $i\notin V(F)$, $\mu^{F}_{i}=\mathtt{DA}_{i}(P)$.
Our analysis builds upon a well-known result which connects improvements over DA to cycles in envy digraphs:
Lemma 1 (Tang and Yu (2014)).
Every Pareto improvement over DA corresponds to a set of trading cycles in $G^{\mathtt{DA}(P)}$. [^3] Similar arguments have been previously made in the literature by Abdulkadiroğlu et al. (2009); Kesten (2010), Erdil (2014) and Ortega et al. (2024).
Having established the required framework, we are now ready to state our results.
4 Limitations of Existing Mechanisms
How many students do existing mechanisms like EADA and DA+TTC actually improve compared to the maximum improvement possible over DA? To quantify this gap, we define the improvement ratio of a mechanism as follows. Let $\mathcal{I}(M,P)$ denote the set of students improved by the mechanism $M$ that Pareto-dominates DA in problem $P$:
$\mathcal{I}(M,P)\coloneqq\{i\in I:\textup{rk}_{i}(M_{i}(P))<\textup{rk}_{i}(\mathtt{DA}_{i}(P))\}$ (2)
We define the maximum improvement $\mathcal{I}(\mu^{*},P)$ as:
$\mathcal{I}(\mu^{*},P)\coloneqq\max_{\mu\in\mathcal{M}^{\mathtt{DA}}(P)}|\mathcal{I}(\mu,P)|$ (3)
Let $\mathcal{P}_{n}$ denote the class of all school choice problems with $n$ students. For any mechanism $M$ that Pareto-dominates DA, we define its improvement ratio as:
$\rho(M;n)\coloneqq\max_{P\in\mathcal{P}_{n}}\frac{|\mathcal{I}(\mu^{*},P)|}{|\mathcal{I}(M,P)|}$ (4)
For example, an improvement ratio of 2 means that the maximum mechanism may generate twice as many improvements as mechanism $M$. Our first result shows that this ratio can grow linearly with the size of the problem. [^4] Throughout Sections 4-4.1, we assume all students consent to waive their priorities, as even with full consent we obtain the worst possible bounds. We examine partial consent explicitly in Section 4.2 and formalize arbitrary consent structures in Section 5.
Theorem 1.
For every $n>4$ there exists a problem $P$ such that the maximum improvement is $n-1$; while that of EADA and DA+TTC are both exactly 2:
$\rho(\text{EADA};n)=\rho(\text{DA+TTC};n)=\frac{n-1}{2}$
Proof.
Upper Bound. First we prove that the improvement ratio cannot exceed $\frac{n-1}{2}$. For the numerator: At most $n-1$ can improve over DA, as there is always at least one unimprovable student (Tang and Yu, 2014). For the denominator: The minimum number of students improved by EADA and DA+TTC is 2, because a trading cycle requires at least two students. [^5] We have assumed that DA is inefficient; otherwise, the minimum would be 0 and the ratio undefined. Therefore, the ratio cannot exceed $\frac{n-1}{2}$.
Tightness. We construct a school choice problem with $n$ students and schools (all with unit capacity) where this improvement ratio is achieved for an arbitrary value of $n>4$.
Let school priorities be such that, for every school $s_{k}$ except $s_{1}$:
$s_{k}:i_{k}\triangleright_{s_{k}}i_{k-1}$ (5)
Students’ preferences are such that, for every student $i_{k}\in I\setminus\{i_{1},i_{n-2},i_{n-1},i_{n}$}:
$i_{k}:s_{1}\succ_{i_{k}}s_{k+1}\succ_{i_{k}}s_{k}$ (6)
The remaining priorities and preferences are as follows:
$s_{1}$ $:i_{1}\triangleright_{s_{1}}i_{n}\triangleright_{s_{1}}i_{2}\triangleright_{s_{1}}i_{n-1}$ (7)
$i_{1}$ $:s_{2}\succ_{i_{1}}s_{n-1}\succ_{i_{1}}s_{1}$ (8)
$i_{n-2}$ $:s_{1}\succ_{i_{n-2}}s_{2}\succ_{i_{n-2}}s_{n-2}$ (9)
$i_{n-1}$ $:s_{1}\succ_{i_{n-1}}s_{2}\succ_{i_{n-1}}s_{n-1}$ (10)
$i_{n}$ $:s_{1}\succ_{i_{n}}s_{n}$ (11)
The table below demonstrates this construction for $n=7$. The DA allocation appears in bold. The allocation that maximises the number of improvements appears in circles, whereas the matching recommended by both EADA and DA+TTC appears in squares (whenever different from DA).
Example 1
$i_{1}$ $i_{2}$ $i_{3}$ $i_{4}$ $i_{5}$ $i_{6}$ $i_{7}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $s_{6}$ $s_{7}$
$s_{2}$ $s_{1}$ $s_{1}$ $s_{1}$ $s_{1}$ $s_{1}$ $s_{1}$ $\bm{i_{1}}$ $\bm{i_{2}}$ $\bm{i_{3}}$ $\bm{i_{4}}$ $\bm{i_{5}}$ $\bm{i_{6}}$ $\bm{i_{7}}$
$s_{6}$ $s_{3}$ $s_{4}$ $s_{5}$ $s_{2}$ $s_{2}$ $\bm{s_{7}}$ $i_{7}$ $i_{1}$ $i_{2}$ $i_{3}$ $i_{4}$ $i_{5}$ $i_{6}$
$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $\bm{s_{5}}$ $\bm{s_{6}}$cdot $i_{2}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
We now demonstrate that the unique stable matching assigns each student $i_{k}$ to the school with the corresponding index $s_{k}$.
Execution of DA.
In DA’s round 1, student $i_{1}$ applies to $s_{2}$, whereas all other students apply to $s_{1}$. School $s_{1}$ temporarily accepts $i_{n}$ (its highest-priority applicant) and rejects all others. In round 2, rejected students $i_{2}$ through $i_{n-1}$ make their next applications. Students $i_{2}$ through $i_{n-3}$ apply to their second-choice schools $s_{3}$ through $s_{n-2}$, while $i_{n-2}$ and $i_{n-1}$ both apply to $s_{2}$. School $s_{2}$ temporarily accepts $i_{1}$ (who applied in round 1) and rejects $i_{n-2}$ and $i_{n-1}$. In subsequent rounds, a cascade of rejections occurs as students apply to their next preferred schools, eventually resulting in each student $i_{k}$ being matched with school $s_{k}$.
The uniqueness of the set of stable matchings can be verified by observing that the school-proposing DA algorithm yields the same outcome, with each school immediately being matched to its highest-priority student.
Analysis of EADA.
Under EADA, we begin by identifying under-demanded schools. First, $s_{n}$ is under-demanded and is removed together with $i_{n}$. In the next iteration, $s_{3},s_{4},\ldots,s_{n-1}$ are under-demanded and removed jointly with $i_{3},i_{4},\ldots,i_{n-1}$, respectively. At this stage, only schools $s_{1}$ and $s_{2}$ remain with students $i_{1}$ and $i_{2}$. These students form a trading cycle where $i_{1}$ improves to $s_{2}$ and $i_{2}$ improves to $s_{1}$. Thus, EADA improves exactly 2 students (assuming $i_{n}$ consents to waive his priority at $s_{1}$, otherwise no student improves).
Analysis of DA+TTC.
In DA+TTC, we begin with the DA allocation where each student $i_{k}$ is assigned to school $s_{k}$. In the TTC algorithm, only students $i_{1}$ and $i_{2}$ form a trading cycle, as $i_{1}$ points to $i_{2}$ and $i_{2}$ points to $i_{1}$. After executing this trade, students $i_{3}$ through $i_{n}$ remain at their DA allocations which are preferred to any remaining allocation. Therefore, DA+TTC also improves exactly 2 students.
Maximum Improvement.
There exists, however, a matching $\mu^{*}$ that dominates DA, obtained by implementing the cycles $(i_{1}\rightarrow i_{n-1}\rightarrow i_{1})$ and $(i_{2}\rightarrow i_{3}\rightarrow\ldots\rightarrow i_{n-2}\rightarrow i_{2})$, which improves $n-1$ students. Under this matching, student $i_{n}$ receives the same allocation as under DA, while all other students receive a more preferred school.
Therefore, the improvement ratio equals $\frac{|I(\mu^{*},P)|}{|I(M,P)|}=\frac{n-1}{2}$ for any $n>4$, establishing the tightness of our bound. ∎
Three remarks are in order now. First, Kesten (2010) already pointed out in his seminal paper that larger improvements than EADA were possible. Our contribution here is to quantify how much larger those improvements can be, leading to an improvement gap that grows linearly with the problem size.
Second, in our constructed example, the maximum improvement also dominates on rank positions improved: the 2-student cycle generates 4 total rank improvements while the $(n-1)$-student cycle generates $n$ total rank improvements. For $n\geq 6$, maximizing student count yields strictly higher rank improvement, showing that existing mechanisms can simultaneously underperform on both distributional and rank-efficiency dimensions, albeit this need not be the case in general.
Third, in our example, the maximum improvement comes at the cost of generating more blocking pairs. In the particular example with seven agents, the maximum improvement generates at least three blocking pairs ($\{i_{1},s_{2}\},\{i_{2},s_{1}\},\{i_{7},s_{1}\}$) versus a single blocking pair generated by EADA and DA+TTC ($\{i_{7},s_{1}\}$). We do not take a stance on whether the larger improvement is worth the additional justified envy, our point is to make explicit that such reduction in justified envy comes at a high cost in terms of how many students are benefited. Furthermore, as we will show next, the larger improvement can sometimes lead to fewer blocking pairs.
4.1 Larger Improvements with Less Justified Envy
Theorem 1 shows that EADA and DA+TTC can improve significantly fewer students, establishing a tight bound. In the example used in the proof, we showed that larger improvements came at a cost of more blocking pairs. Now we will show that this need not always be the case: in some problems, EADA and DA+TTC can be doubly dominated: they may improve fewer students while generating more justified envy.
Proposition 1.
There exist school choice problems where a larger improvement than EADA’s and DA+TTC’s is possible that generates fewer blocking pairs.
Proof.
Comparison with EADA. Consider the following school choice problem with seven students and seven schools with one seat each. DA’s allocation appears in bold. Preferences and priorities left unspecified are irrelevant.
Example 2
$i_{1}$ $i_{2}$ $i_{3}$ $i_{4}$ $i_{5}$ $i_{6}$ $i_{7}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $s_{6}$ $s_{7}$
$s_{6}$ $s_{1}$ $s_{6}$ $s_{5}$ $s_{3}$ $s_{4}$ $s_{4}$ $\bm{i_{1}}$ $\bm{i_{2}}$ $\bm{i_{3}}$ $\bm{i_{4}}$ $\bm{i_{5}}$ $\bm{i_{6}}$cdot
$s_{4}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $s_{6}$ $\boldsymbol{s_{6}}$ $\boldsymbol{s_{7}}$ $i_{5}$ $i_{1}$ $i_{5}$ $i_{7}$ $i_{4}$ $i_{3}$cdot
$s_{2}$cdot $\cdot$ $\cdot$ $s_{4}$cdot $\cdot$ $i_{2}$cdot $i_{1}$ $i_{1}$ $i_{1}$ $i_{5}$cdot
$s_{3}$cdot $\cdot$ $\cdot$ $s_{1}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$cdot $i_{1}$cdot
$s_{5}$cdot $\cdot$ $\cdot$ $\bm{s_{5}}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{5}$cdot $\cdot$ $\cdot$
$\bm{s_{1}}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
The DA envy digraph appears below, showing the trading cycles implemented by EADA (Figure 1, improving 4 students through the cycle $i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$) and the maximum improvement (Figure 2, improving 6 students via two disjoint cycles, namely $i_{1}\to i_{2}\to i_{1}$ and $i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3}$). Interestingly, EADA generates 3 blocking pairs ($\{i_{3},s_{6}\},\{i_{5},s_{6}\},\{i_{7},s_{4}\}$), whereas the maximum improvement generates only two blocking pairs ($\{i_{1},s_{4}\},\{i_{7},s_{4}\}$). [^6] Note that here we are simply counting the number of blocking pairs; there is a literature (see Section 2) arguing that that some are more relevant than others. Note that the maximum improvement improves every improvable student, because $i_{7}$ is unimprovable since nobody envies him.
$i_{1}$$i_{2}$$i_{3}$$i_{4}$$i_{5}$$i_{6}$$i_{7}$$i_{1}$$i_{4}$$i_{5}$$i_{6}$
Figure 1: EADA
$i_{1}$$i_{2}$$i_{3}$$i_{4}$$i_{5}$$i_{6}$$i_{7}$$i_{1}$$i_{2}$$i_{3}$$i_{4}$$i_{5}$$i_{6}$
Figure 2: Maximum Improvement
Comparison with DA+TTC. To show that a larger improvement can generate fewer blocking pairs than DA+TTC and can be implemented with fewer students consenting, consider the following example with five students and five schools with unit capacity. DA’s allocation appears in bold.
Example 3
$i_{1}$ $i_{2}$ $i_{3}$ $i_{4}$ $i_{5}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$
$s_{2}$ $s_{1}$ $s_{1}$ $s_{1}$ $s_{1}$ $\bm{i_{1}}$ $\bm{i_{2}}$ $\bm{i_{3}}$ $\bm{i_{4}}$ $\bm{i_{5}}$
$s_{4}$ $s_{3}$ $s_{2}$ $s_{2}$ $\bm{s_{5}}$ $i_{5}$ $i_{4}$ $i_{4}$ $i_{5}$cdot
$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$cdot $i_{4}$ $i_{3}$ $i_{2}$ $i_{1}$cdot
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{2}$ $i_{1}$cdot $\cdot$ $\cdot$
In DA+TTC, students $i_{1}$ and $i_{2}$ trade their DA allocations, generating four blocking pairs: $(i_{4},s_{1}),(i_{5},s_{1}),(i_{3},s_{2}),(i_{4},s_{2})$. In contrast, the maximum improvement is the cycle $(i_{1}\to i_{4}\to i_{1}),(i_{2}\to i_{3}\to i_{2})$, generating only one blocking pair: $(i_{5},s_{1})$. ∎
While DA+TTC, and in particular, EADA, choose improvement cycles taking priorities into account, therefore satisfying several weak stability properties (see Section 2), Proposition 1 shows that they may generate doubly dominated matchings in some problems—improving fewer students while creating more blocking pairs—since alternative matchings can achieve better outcomes on both dimensions simultaneously.
Noting that the larger improvements in the previous two examples are obtained by DA+TTC and EADA, respectively, yields the following incomparability corollary.
Corollary 1.
EADA and DA+TTC are incomparable in terms of size of improvement and number of blocking pairs, i.e. there exist problems where DA+TTC doubly dominates EADA and vice versa.
We find Corollary 1 interesting because DA+TTC has been considered suboptimal regarding a number of axioms (Kesten, 2010; Troyan et al., 2020), yet it is not uniformly doubly dominated by the more well-regarded EADA mechanism.
4.2 Consent and Improvement Size
In his seminal paper, Kesten (2010) argues that “the likelihood of getting stuck without gaining much Pareto improvement clearly increases with the cycle size". This intuition seems natural—coordinating more students should be harder. However, this relationship is more nuanced than originally thought.
Let us recall that, in EADA, a student agrees to waive his priorities before the DA mechanism is executed at schools where he is an interrupter. Kesten shows that such students never benefit nor gain from giving their consent. Let us divide students who consent into two classes: non-beneficiary consenters are those who neither benefit nor are harmed by giving their consent, and beneficiary consenters who explicitly benefit from the trading cycles generated by their consent to waive their priorities. It is reasonable to conjecture that beneficiary consents are more likely to be obtained than non-beneficiary ones, as the former ones directly benefit rather than providing consent altruistically.
With this distinction in mind, let us revisit Example 2 in the proof of Proposition 1. EADA implements the four-cycle ($i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$), requiring non-beneficiary consent from both $i_{3}$ (at $s_{6}$) and $i_{7}$ (at $s_{4}$), as well as beneficiary consent from $i_{5}$ (at $s_{6}$). In contrast, the maximum improvement implements two disjoint cycles requiring fewer non-beneficiary consents (1 versus 2 of EADA), and the same number of beneficiary consents (one each, see Table 1). If non-beneficiary consents, which rely purely on altruism, fail, EADA cannot improve any student, and yet the largest improvement can still execute the two-cycle ($i_{1},i_{2},i_{1}$), given the consent of potential beneficiary $i_{5}$. Student $i_{5}$ does not benefit given the lack of non-beneficiary consents, but since consents are expressed ex-ante this does not affect his decision on whether to consent or not.
Table 1: Consent Requirements in Example 2
Mechanism Consent Required Improvement Size
Non-beneficiary Beneficiary
EADA $i_{3},i_{7}$ $i_{5}$ 4 ($i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$)
$i_{4}$, $i_{3}$ (no $i_{7}$) 2 ($i_{1}\to i_{5}\to i_{1}$)
$i_{5}$ (no $i_{7},i_{4}$) 2 ($i_{1}\to i_{2}\to i_{1}$)
otherwise 0
Maximum $i_{7}$ $i_{1}$ 6 ($i_{1}\to i_{2}\to i_{1}$), ($i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3}$)
Improvement $i_{5}$ 2 ($i_{1}\to i_{2}\to i_{1}$)
otherwise 0
We summarize this observation as Proposition 2 below.
Proposition 2.
There exist problems for which a Pareto improvement that strictly improves more students than EADA requires fewer total and non-beneficiary consents.
Furthermore, in some problems, larger improvements can be implemented relying solely on beneficiary consents, whereas EADA may fail to implement smaller improvements without non-beneficiary consent.
Having examined the potential shortcomings of existing mechanisms, we proceed to analyse a new mechanism that improves as many students as possible compared to the DA baseline.
5 Maximum Implementable Improvement over Deferred Acceptance
Our theoretical analysis has revealed that existing mechanisms can fall short of maximizing the number of students who benefit from efficiency enhancements to DA. The Maximum Implementable Improvement over Deferred Acceptance (MIIDA) mechanism addresses this limitation by identifying and implementing improvements that benefit the maximum possible number of students given the consent structure.
To define MIIDA formally, let us incorporate consent into the mathematical framework. Each student reports strict preferences over schools $\succ_{i}$ and whether they consent $c_{i}=1$ (or not, with $c_{i}=0$) to waive their priority at schools. A consent-based mechanism maps a school choice problem and a consent decision by each student to a matching. Let $C\subseteq I$ denote the set of students who agree to waive their priority when they do not benefit from the proposed trades. For any edge $i\to j$ in the envy digraph $G^{\mathtt{DA}(P)}$, let $\text{consent}(i\to j)$ denote the set of students whose consent is required for student $i$ to receive student $j$’s DA assignment. Specifically,
$\text{consent}(i\to j)=\{k\in I:k\text{ has higher priority than }i\text{ at school }\mathtt{DA}_{j}(P)\}$
For any feedback set $F$, we define $L(F)=\bigcup_{C\in F}\bigcup_{(i\to j)\in C}\text{consent}(i\to j)$ as the set of all students whose consent is required to implement the trades in $F$. A feedback set $F$ is implementable given consent set $C$ if $L(F)\setminus V(F)\subseteq C$, so that all required consents come from either beneficiaries of the trades or students in the consent set $C$.
We define MIIDA as follows.
Algorithm 1 Maximum Improvement over Deferred Acceptance (MIIDA)
1:Compute the student-optimal stable matching $\mu^{0}=\mathtt{DA}(P)$
2:Set iteration counter $t=0$
3:repeat
4:     Generate the envy digraph $G^{\mu^{t}}$ for current matching $\mu^{t}$
5:     if $G^{\mu^{t}}$ contains no cycles then
6:         break
7:     end if
8:     Let $N^{t}=I\setminus C$ be the set of non-consenting students
9:     Remove nodes $N^{t}$ and adjacent edges from $G^{\mu^{t}}$ to obtain reduced graph $\tilde{G}^{\mu^{t}}$
10:     if $\tilde{G}^{\mu^{t}}$ contains no cycles then
11:         break $\triangleright$ No improvements possible with available consent
12:     end if
13:     Find maximum feedback set $F^{t}$ in $\tilde{G}^{\mu^{t}}$ (with a fixed tie-breaking rule)
14:     Implement trades in $F^{t}$ to produce matching $\mu^{t+1}$
15:     $t\leftarrow t+1$
16:until convergence
17:return final matching $\mu^{t}$
MIIDA iteratively finds maximum implementable feedback sets in the envy digraph until no further improvements are possible. Iteration is essential because implementing trades changes the envy structure: after students exchange assignments, new envy relationships emerge while others disappear. To illustrate, consider the consent-labeled envy digraph from our earlier example. MIIDA proceeds as follows:
  1. 1.
    In the first iteration, MIIDA identifies the maximum feedback set
    $F={(i_{1}\to i_{2}\to i_{1}),(i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3})}$
    covering 6 students, which requires consent from $L(F)\setminus V(F)={i_{7}}$.
  2. 2.
    If $\{i_{1},i_{7}\}\subset C$, MIIDA implements $F$.
  3. 3.
    If the required consents are not available, MIIDA seeks the largest implementable feedback set in the graph. If $i_{7}\in C$, it finds $F^{\prime}={(i_{1}\to i_{2}\to i_{1})}$ covering 2 students.
  4. 4.
    If $i_{5}\in C$, implement $F^{\prime}$. Otherwise, make no changes over DA.
MIIDA satisfies a number of important desiderata.
Theorem 2.
MIIDA has the following properties:
  1. 1.
    Weakly Pareto-dominates DA.
  2. 2.
    Is Pareto-efficient if all students consent.
  3. 3.
    Only violates the priorities of students who consent.
  4. 4.
    Maximises the number of students who benefit from Pareto improvements over DA given the consent structure (so it is never doubly-dominated).
  5. 5.
    Can generate strictly fewer, but also strictly more, blocking pairs than EADA and DA+TTC, depending on the school choice problem.
  6. 6.
    No student has incentives not to consent. Specifically, for every student consenting is a very weakly dominant strategy in the complete information preference revelation game induced by MIIDA in which students simultaneously report their preferences and whether they consent to waive their priorities. [^7] A strategy is very weakly dominant for a player if, for every strategy profile, it gives a weakly higher utility than every other strategy (e.g. Chen and Micali, 2015). On the other hand, a weakly dominant strategy additionally requires that for some strategy profile it must yield a strictly higher utility than any other strategy available to the player.
  7. 7.
    Can be computed in polynomial time.
Proof.
  1. 1.
    Pareto-dominance: Follows directly from MIIDA’s construction: it implements only cycles where each participant receives a strictly preferred school relative to DA (by envy digraph definition). Non-participants keep their DA assignments, guaranteeing weak Pareto dominance for everyone.
  2. 2.
    Pareto-efficient if all consent: When $C=I$, $\tilde{G}^{\mu^{t}}=G^{\mu^{t}}$. Termination occurs only when no cycles exist, implying no further Pareto improvements are possible (by Lemma 1)).
  3. 3.
    Priority violations require consent: For any edge $i\to j$ implemented, consent is required from all $k$ with higher priority at $s=\mathtt{DA}_{j}(P)$. The condition $L(F^{t})\setminus V(F^{t})\subseteq C$ ensures only consenting non-beneficiaries’ priorities are violated.
  4. 4.
    Maximizes improvements given consent: Let $S$ be students improved by MIIDA. Suppose another mechanism $M$ improves $T$ with $|T|>|S|$. By Lemma 1, $T$ corresponds to disjoint cycles $\mathcal{C}_{T}$ in $G^{\mathtt{DA}(P)}$ implementable under $C$. In MIIDA’s first iteration, it selects a feedback set $F^{0}$ maximizing $|V(F^{0})|$ over implementable sets. Since $\mathcal{C}_{T}$ is implementable: $|V(F^{0})|\geq|T|$. But $S$ includes $V(F^{0})$ (at minimum), so $|S|\geq|T|$, contradiction. Thus no such $M$ exists.
  5. 5.
    Variable blocking pairs: Proven by Theorem 1 and Proposition 1: Example 1 shows MIIDA can generate more blocking pairs than EADA/DA+TTC; Example 2 shows it can also generate fewer.
  6. 6.
    Consent incentives: Fix student $i\in I$, any preference profile $P=(\succ^{\prime}_{i},\succ^{\prime}_{-i})$, and any consent decisions $c_{-i}$ by other students. Let $\mu^{\mathtt{DA}(P)}=\mathtt{DA}(P)$ and let $C=\{j\neq i:c_{j}=1\}\cup\{i\}$ denote the consent set when student $i$ consents. Let $F$ denote the maximum implementable feedback set under consent set $C$.
    When $c_{i}=0$, by MIIDA’s construction, $i$ cannot participate in any trading cycle. Therefore $\mathtt{MIIDA}_{i}(P;c_{i}=0,c_{-i})=\mu^{\mathtt{DA}(P)}_{i}$.
    When $c_{i}=1$:
    • If $i\notin V(F)=\mathtt{MIIDA}_{i}(P;c_{i}=1,c_{-i})=\mu^{\mathtt{DA}(P)}_{i}$
    • If $i\in V(F)=\mathtt{MIIDA}_{i}(P;c_{i}=1,c_{-i})=\mu^{\mathtt{DA}(P)}_{\text{succ}(i)}$
    where $\text{succ}(i)$ is $i$’s successor in the cycle containing $i$ within $F$. If $i\in V(F)$, then by definition of envy digraph $G^{\mu^{\mathtt{DA}(P)}}$, there exists edge $i\to\text{succ}(i)$, which means $\mu^{\mathtt{DA}(P)}_{\text{succ}(i)}\succ^{\prime}_{i}\mu^{\mathtt{DA}(P)}_{i}$. Therefore $\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{\text{succ}(i)})<\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{i})$.
    It follows that
    $\textup{rk}_{i}(\mathtt{MIIDA}(P;c_{i}=1,c_{-i}))\leq\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{i})=\textup{rk}_{i}(\mathtt{MIIDA}(P;c_{i}=0,c_{-i}))$
    with strict inequality when $i\in V(F)$. Since this holds for arbitrary $(\succ^{\prime}_{i},\succ^{\prime}_{-i},c_{-i})$, consenting is a very weakly dominant strategy for student $i$.
  7. 7.
    Polynomial runtime: MIIDA runs in polynomial time through its iterative structure. DA is computed once initially in $O(n^{2})$ time (Gale and Shapley, 1962). Each subsequent iteration involves: (i) generating the envy digraph for the current matching in $O(n^{2})$ time, and (ii) finding a maximum feedback set. The cycle packing problem can be solved in $O(m\sqrt{n})$ time where $m$ is the number of edges and $n$ is the number of vertices (Biro et al., 2009). Since the envy digraph has at most $n^{2}$ edges, this step takes $O(n^{2}\sqrt{n})=O(n^{2.5})$ time per iteration. Since each iteration covers at least one additional student in trading cycles, and at most $n-1$ students can be improved over DA (Tang and Yu, 2014), MIIDA terminates within $n$ iterations. Therefore, the overall complexity is $O(n^{3.5})$.
While MIIDA improves as many students as possible given the consent structure available, and is never doubly dominated, it also has a number of important limitations. Arguably, the most severe is its lack of stability. Since it chooses cycles without taking into account priorities, it can lead to more blocking pairs than EADA and DA+TTC (but also fewer, as we have seen). In fact, MIIDA sometimes generates a set of blocking pairs that is a strict superset of EADA’s blocking pairs. Unfortunately, this is unavoidable by any mechanism that does not miss a significant number of improvement cycles.
To state this result formally, let us denote by $B(M,P)$ the set of blocking pairs generated by mechanism $M$ in problem $P$. We say mechanism $M$ is setwise justified envy minimal if for any problem $P$, there is no alternative mechanism $M^{\prime}$ such that $B(M^{\prime},P)\subsetneq B(M,P)$. A number of mechanisms, including EADA (Doğan and Ehlers, 2021a), satisfy this partial comparison property. However, setwise justified envy minimality is at odds with larger improvements.
Theorem 3.
Among Pareto-efficient mechanisms that weakly Pareto-dominate DA, every mechanism that is setwise justified envy minimal achieves an improvement ratio of $\frac{n-1}{2}$.
Proof.
Recall Example 1 in the proof of Theorem 1. In the case with $7$ students and schools, there are six cycles in DA’s envy digraph, summarized in Table 2.
Table 2: Cycles and Blocking Pairs in Example 1
Cycles Blocking Pairs
$K_{1}=(i_{1}\to i_{2}\to i_{1})$ $(i_{7},s_{1})$
$K_{2}=(i_{1}\to i_{2}\to i_{3}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$K_{3}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$K_{4}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$K_{5}=(i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2})$ $(i_{1},s_{2})$
$K_{6}=(i_{1}\to i_{6}\to i_{1})$ $(i_{7},s_{1}),(i_{2},s_{1}),(i_{5},s_{6})$
$K_{7}=(i_{1}\to i_{6}\to i_{2}\to i_{1})$ $(i_{7},s_{1}),(i_{1},s_{2}),(i_{5},s_{6})$
Any efficient mechanism that weakly dominates DA must implement at least one of these cycles (potentially two if mutually disjoint). In particular, MIIDA implements ($i_{1}\to i_{6}\to i_{1}$) and ($i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2}$).
The cycles $K_{2},K_{3},K_{4},K_{6},K_{7}$ all generate a set of blocking pairs that is strictly larger (by set inclusion) than that generated by $K_{1}$. The cycle $K_{5}$ is incomparable to $K_{1}$, but implementing it by itself does not lead to a Pareto-efficient allocation, since the cycle $K_{6}$ could be jointly implemented. But if both $K_{5}$ and $K_{6}$ are implemented, the set of blocking pairs becomes $\{(i_{1},s_{2}),(i_{2},s_{1}),(i_{7},s_{1})\}$, a strict superset of the set of blocking pairs generated by $K_{1}$. Thus, we conclude that every setwise justified envy minimal mechanism must implement $K_{1}$therefore missing the larger improvements dictated by $K_{5}$ and $K_{6}$ jointly, leading to an improvement ratio of $\frac{n-1}{2}=\frac{6}{2}$.
Finally, the extension to arbitrary $n$ is straightforward: there are $n$ cycles, with $K_{k}=(i_{1}\to\ldots\to i_{k+1}\to i_{1})$ for all integer $31\leq k\leq n-$ 3, all of them generating the blocking pair ($i_{n},s_{1}$), and only $K_{1}$ introducing exactly this blocking pair. There are exactly three more cycles: $K_{n-2}=(i_{2}\to i_{3}\to\ldots\to i_{n-3}\to i_{2})$, $K_{n-1}=(i_{1}\to i_{6}\to i_{1})$, and $K_{n}=(i_{1}\to i_{n-1}\to i_{2}\to i_{1})$. Only one collection of cycles that leaves the graph acyclic is setwise justifiable envy minimal: $K_{1}$. However, the cycles $K_{n-2}$ and $K_{n-1}$ are compatible and lead to an allocation where $n-1$ students improve, and therefore any setwise justified envy minimal mechanism incurs in the largest improvement ratio possible, namely $\frac{n-1}{2}$. ∎
MIIDA also has important limitations beyond its lack of stability properties. The mechanism is manipulable and may have multiple solutions. Nonetheless, despite its limitations, we believe MIIDA serves as a useful mechanism to help us better understand the tradeoffs between efficiency and stability when improving over the celebrated DA mechanism.
6 Conclusion
We have shown that EADA and DA+TTC can be doubly dominated, improving as few as 2 students when $n-1$ could benefit, while generating more blocking pairs than larger improvements. MIIDA addresses this by maximizing the number of students who benefit, though our impossibility theorem proves this necessarily comes at the cost of losing EADA’s minimal justified envy guarantee.
An equally promising alternative would modify how EADA handles consent to minimize blocking pairs rather than maximize improvements. While such a mechanism might not improve as many students as MIIDA, it could nonetheless avoid double-dominance. This represents a complementary approach to navigating the efficiency-stability frontier that warrants future investigation.

### v3

@@
-Maximum Implementable Improvements over Deferred Acceptance
+The Trade-off Between Minimal Instability and Larger Improvements over Deferred Acceptance
 Abstract
-The Deferred Acceptance (DA) mechanism produces stable but Pareto-inefficient school assignments.
-While mechanisms like Efficiency-Adjusted DA address this inefficiency by asking students to waive priorities at unattainable schools, they can leave significant gains unrealized—sometimes improving only two students when all but one could benefit.
-Surprisingly, such mechanisms can be doubly dominated, creating fewer improvements while generating more justified envy.
-We introduce the Maximum Implementable Improvement over the DA mechanism (MIIDA), which maximizes the number of students made better off given priority-waiving consents.
-MIIDA weakly Pareto dominates DA, is never doubly dominated, and incentivizes students to consent.
-school choice,
-maximum improvements,
-consent,
-C78,
+The celebrated Efficiency-Adjusted Deferred Acceptance mechanism (EADA) improves the efficiency of the DA algorithm via consented priority violations.
+Notwithstanding its many merits, we show that EADA can improve only two students when an alternative mechanism that Pareto-dominates DA could benefit all but one student.
+This shortfall in the number of students improved is not exclusive of EADA but extends to all setwise minimally unstable mechanisms, i.e.
+those that generate a set of blocking pairs that is never a strict superset of that of another mechanism.
+The incompatibility between number of students improved and minimal instability disappears when blocking pairs are compared cardinally rather than by set inclusion.
+In some problems, EADA can be doubly dominated: improving fewer students while generating more blocking pairs.
 D47,
@@
 […]
+[…]
 Knipe: tknipe01@qub.ac.uk, Ortega: j.ortega@qub.ac.uk.
-We acknowledge helpful discussions with Pablo Arribillaga, Péter Biró, Battal Doğan, Martin Meier, and Gabriel Ziegler.
+We acknowledge helpful discussions with Péter Biró and Battal Doğan.
 T.
@@
 1 Introduction
-The student-proposing Deferred Acceptance (DA) algorithm has become a pillar of market design, combining elegant theoretical properties with widespread adoption in school assignment systems worldwide.
-While DA guarantees stable matchings, it can produce allocations that are Pareto-inefficient for students in theory and in practice.
-This inefficiency has motivated the development of mechanisms that weakly Pareto-dominate DA by implementing mutually beneficial trades among students, which only violate the priorities of students who willingly consent to such violations.
-The most prominent solutions—Efficiency-Adjusted DA (EADA, Kesten 2010) and Top Trading Cycles applied to DA assignments (DA+TTC)—fix DA’s inefficiency without harming any student while, at the same time, preserving weak but meaningful stability and incentive properties.
-In particular, EADA has shown a remarkable performance in both theory and the lab, and has been considered for practical implementation in Flanders (discussed in Section 2).
-In this paper, we analyze how many students can improve their DA assignment through the trading cycles implemented by EADA and DA+TTC, while tracking the number of blocking pairs these exchanges introduce.
-We find that both mechanisms can perform poorly when judged solely by the number of students they improve.
-In problems with $n$ students, we can construct examples where $n-1$ students could benefit.
-However, EADA and DA+TTC forgo this large cycle involving almost all students and perform a single two-way exchange instead.
-Therefore, the improvement ratio of both mechanisms, which tells us how many efficiency swaps they miss, is $\frac{n-1}{2}$ (the largest possible ratio obtained by an efficient mechanism that improves upon DA) and grows linearly with the size of the school choice problem (Theorem 1).
-Surprisingly, we find that existing mechanisms can be doubly dominated, simultaneously improving fewer students while generating more blocking pairs in specific problems than alternative allocations (Proposition 1).
-This counterintuitive result reveals an unexploited margin on the efficiency-stability frontier: some Pareto improvements strictly dominate EADA and DA+TTC on both dimensions, yet mechanisms that achieve this have not been explored in the literature.
-These larger improvements can sometimes require fewer consents overall, and specifically fewer non-beneficiary consents—agreements from students who do not directly benefit from the proposed trades and must therefore act altruistically—compared to the smaller improvements implemented by existing mechanisms (Proposition 2).
-Furthermore, EADA and DA+TTC prove incomparable regarding double dominance—in different problem instances, each can produce matchings that doubly dominate the other.
-This mutual non-dominance suggests that the choice between EADA and DA+TTC involves case-specific tradeoffs rather than universal superiority.
-The potentially small number of improvements carried out by existing mechanisms motivates us to propose a new mechanism that always improves the placement of as many students as possible.
-This algorithm computes DA and, based on the corresponding envy digraph, chooses the collection of disjoint trading cycles that maximizes the number of students involved given the consent structure available.
-We call this mechanism Maximum Implementable Improvement over Deferred Acceptance (MIIDA).
-MIIDA weakly Pareto dominates DA and is Pareto-efficient when all students consent.
-It can generate less or more justified envy, depending on the problem, but it is never doubly dominated because it improves as many students as possible.
-With regard to incentives, it gives every student a weak incentive to consent to waive their priorities.
-Finally, it can be computed in polynomial time (Theorem 2).
-While MIIDA provides valuable theoretical insights about mechanism performance, it has important drawbacks too: it can be manipulated and can generate more blocking pairs than EADA and DA+TTC.
-Moreover, the set of blocking pairs that it generates may be a strict superset of the blocking pairs generated by other setwise justified envy minimal mechanisms such as EADA.
-Such negative properties are unavoidable: any efficient mechanism that weakly Pareto dominates DA and is setwise justified envy minimal has an improvement ratio $\frac{n-1}{2}$, leaving large efficiency gains unrealized (Theorem 3).
-Our findings show that consent structures already embedded in existing mechanisms contain unexploited potential for achieving greater welfare improvements, as larger improvements can sometimes be achieved while requiring fewer consents overall, and particularly fewer altruistic consents from non-beneficiary students.
-Moreover, these larger improvements can occasionally generate fewer blocking pairs than smaller ones, creating a counterintuitive bidimensional improvement on the efficiency-stability trade-off.
-Overall, MIIDA serves as a useful benchmark that allows us to quantify precisely how much improvement potential is foregone by existing mechanisms and reveals opportunities for further mechanism design in this space.
-To summarize, our findings contribute to the literature in three ways.
-First, we provide tight bounds on the improvement ratios of existing mechanisms, showing that both EADA and DA+TTC can leave significant efficiency gains unrealized in worst-case scenarios, and show that these efficiency gains may be compatible with parallel stability gains.
-Second, we introduce MIIDA as a new mechanism in the relatively under-explored class of efficient mechanisms that Pareto-dominate DA and document its properties and weaknesses.
-Third, we establish an impossibility result: any mechanism that attains a non negligible improvement size cannot be justified envy minimal.
+The student-proposing Deferred Acceptance (DA) algorithm has become a cornerstone of market design, combining elegant theoretical properties with widespread adoption.
+However, while DA generates stable matchings, it can produce allocations that are Pareto-inefficient for students.
+The large size of this inefficiency in practice has motivated the development of mechanisms that Pareto-dominate DA by implementing mutually beneficial trades among students, violating only the priorities of students who willingly consent to such violations.111[Abdulkadiroğlu et al.
+(2009)] report that 4,296 students in NYC would be assigned to a better school without harming any other pupil.
+The most prominent among these is the Efficiency-Adjusted Deferred Acceptance mechanism (EADA, Kesten 2010), which addresses DA’s inefficiency while preserving weak but meaningful stability and incentive properties.
+A large literature has shown that EADA performs remarkably well both theoretically and in the lab, making it a key mechanism in the market design literature and leading to its consideration for practical implementation (see Section 2).
+In this paper, we study how many students EADA actually improves compared to other mechanisms that also Pareto-dominate DA, and whether these mechanisms improve more students while also generating less instability (as any mechanism that Pareto-dominates DA is bound to create some blocking pairs).
+Despite EADA’s excellent performance with regard to other criteria, we find that it can perform poorly when judged solely by the number of students it improves.
+In school choice problems with $n$ students, we construct examples where n1n-1 students could benefit and yet EADA executes only a single two-way exchange instead (even when all students consent to waive their priorities).
+Therefore, EADA’s improvement ratio, defined as the size of the largest improvement possible divided by the size of the actual improvement implemented in the worst-case instance, is $\frac{n-1}{2}$, which is the largest possible ratio obtained by any Pareto-efficient mechanism that improves upon DA and grows linearly with $n$ (Theorem 1).
+Moreover, the concern that only a few students may improve upon their DA placement is not unique to EADA.
+In fact, any setwise minimally unstable mechanism that Pareto-dominates DA shares the same limitation (Theorem 2).
+The impossibility of improving more students while minimizing justified envy in a setwise sense is particularly relevant because each blocking pair represents an unfairly treated student who could potentially challenge the allocation legally.
+Therefore, mechanism designers who insist on setwise minimizing envy while improving upon DA must accept that the distribution of improvements may be rather limited compared to what other mechanisms can achieve.
+Conversely, policymakers who intend to maximize the coverage of improvement across the student population must recognize that they may create a strictly setwise larger number of students who are treated unfairly and thus can challenge the assignment.
+Taken together, Theorems 1 and 2 extend Kesten’s observation that EADA may not improve as many students as possible by establishing exact bounds on this shortfall and proving an impossibility theorem that highlights the underlying tradeoff between improving more students and being setwise minimally unstable, a property satisfied by EADA among other mechanisms (Doğan and Ehlers, 2021a; Tang and Zhang, 2021).
+However, if we compare the instability generated simply in terms of the number of blocking pairs created, we obtain room for improvement in both number of students improved and number of blocking pairs created.
+We show that EADA can be doubly dominated—simultaneously improving fewer students and generating more blocking pairs than alternative mechanisms that also Pareto-dominate DA (Theorem 3).
+This previously unknown bi-dimensional possibility of improvement opens an avenue of research into the relatively under-explored class of mechanisms that Pareto-dominate the ubiquitously used Deferred Acceptance mechanism.
 Outline.
-The remainder of this paper is organised as follows.
-Section 2 discusses the literature.
-Section 3 introduces the model.
-Section 4 presents the limitations of existing mechanisms.
-Section 5 introduces the Maximum Implementable Improvement over Deferred Acceptance mechanism.
-Section 6 concludes.
+The remainder of this paper is organised as follows. Section 2 discusses the literature. Section 3 introduces the model. Section 4 presents our results. Section 5 concludes.
 2 Related Literature
@@
 The Pareto-inefficiency of the student-proposing Deferred Acceptance algorithm has been recognized since the early studies on matching markets and school choice (Roth, 1982; Abdulkadiroğlu and Sönmez, 2003).
-Kesten (2010) provided the first theoretical bounds on DA’s inefficiency, establishing that the welfare losses can be substantial.
-DA’s inefficiency has been consistently documented in practice.
-Studies including Abdulkadiroglu et al.
-(2005), Abdulkadiroğlu et al.
-(2009), Che and Tercieux (2019), Abdulkadiroğlu et al.
-(2020), and Ortega and Klein (2023) consistently find that DA leaves significant welfare gains on the table across different school choice systems.
-On the other hand, theoretical literature has identified restrictive conditions on preference and priority structures under which DA achieves efficiency (e.g.
-Ergin, 2002; Cantillon et al., 2024).
+It has been consistently documented in practice (Abdulkadiroglu et al., 2005; Abdulkadiroğlu et al., 2009; Che and Tercieux, 2019; Abdulkadiroğlu et al., 2020; Ortega and Klein, 2023).
 Mechanisms That Improve Upon DA.
@@
 EADA identifies Pareto-improving trading cycles among students that only violate the priorities of students who benefit from consenting to waive their priorities, ensuring such students have no incentive to withhold consent.
-While Kesten notes that EADA’s improvements may not be the largest possible and argues that selecting larger cycles could be problematic due to the greater number of consents required, our contribution is to quantify exactly how many improvements EADA may be missing and to show that such larger improvements need not always require more consent nor generate more justified envy.
-EADA has been extensively studied both theoretically and experimentally, with the literature consistently documenting its strong performance across multiple stability and incentives criteria (Bando, 2014; Tang and Yu, 2014; Dur et al., 2019; Troyan et al., 2020; Troyan and Morrill, 2020; Ehlers and Morrill, 2020; Tang and Zhang, 2021; Reny, 2022; Chen and Möller, 2023; Cerrone et al., 2024; Cheng et al., 2024; Shirakawa, 2025).
-Form this literature, the papers closer to ours are Doğan and Ehlers (2021a) and Kwon and Shorrer (2020).
-These papers show that EADA (among other mechanisms) is justified envy minimal among efficient mechanisms in a set inclusion sense using blocking pairs or triplets, respectively.
-[^1] Abdulkadiroğlu et al.
-(2020) and Doğan and Ehlers (2021b) show that the Top Trading Cycles algorithm is also (setwise) justified envy minimal.
-While this property has made EADA attractive for practical implementation, we prove that any mechanism satisfying setwise justified envy minimality must have an improvement ratio of at least $\frac{n-1}{2}$, leaving large welfare gains unrealized in specific problems.
-Our MIIDA mechanism deliberately relaxes this constraint to maximize the number of students who benefit, demonstrating that the choice between minimal justified envy and maximal improvement size represents a stark trade-off.
-Another natural approach to improving DA is to apply the Top Trading Cycles mechanism to DA’s allocation.
-While this mechanism is know in the literature, it has received surprisingly limited formal analysis (with some exceptions, see Alcalde and Romero-Medina 2017; Kesten 2010; Troyan et al.
+EADA has been extensively studied both theoretically and experimentally, with the literature consistently documenting its strong performance across multiple stability and incentives criteria.
+From this literature, the papers closest to ours are Doğan and Ehlers (2021a); Tang and Zhang (2021) and Kwon and Shorrer (2020).
+These three papers show that EADA is setwise minimally unstable when using blocking pairs or triplets for the comparison, respectively.
+[^2] Outside of the class of mechanisms that Pareto-dominate DA, Abdulkadiroğlu et al.
+(2020) and Doğan and Ehlers (2021b) show that the Top Trading Cycles algorithm is also setwise minimally unstable.
+We contribute to this literature by showing that any setwise minimally unstable mechanism is bound to improve only a few students when many could alternatively benefit.
+Furthermore, Doğan and Ehlers (2021a) also show that when comparing minimal instability in a cardinal sense, there is no Pareto-efficient mechanism that Pareto-dominates DA that is cardinally minimally unstable.
+In this direction, we further show that it is possible to generate fewer blocking pairs than EADA while improving more students.
+Beyond these papers, the broader literature has consistently praised EADA’s performance across other dimensions (Bando, 2014; Tang and Yu, 2014; Dur et al., 2019; Troyan et al., 2020; Troyan and Morrill, 2020; Ehlers and Morrill, 2020; Tang and Zhang, 2021; Reny, 2022; Chen and Möller, 2023; Cerrone et al., 2024; Cheng et al., 2024; Dogan and Ehlers, 2023; Shirakawa, 2025).
+Another natural approach to Pareto-improve DA is to apply the Top Trading Cycles mechanism to DA’s allocation.
+While this mechanism is known in the literature, it has received surprisingly limited formal analysis (with some exceptions, see Alcalde and Romero-Medina 2017; Kesten 2010; Troyan et al.
 2020).
-The broader literature has explored various other approaches to improving DA.
-Kesten and Kurino (2019) examine conditions under which DA can be improved while maintaining strategy-proofness, while Doğan and Yenmez (2020) study consistency properties of DA improvements.
-Dur and Morrill (2020) examine the equilibrium performance of mechanisms that dominate DA.
-Ortega et al.
-(2025) demonstrate that certain limitations of DA—including rank-inefficiency, inequality and segregation—cannot be addressed by any Pareto-improvement.
+All of our EADA results extend to DA+TTC: it can miss significantly larger improvements while generating more blocking pairs.
+More generally, a series of papers have examined when it is possible to improve upon DA while maintaining strategy-proofness (Kesten and Kurino, 2019) or consistency (Doğan and Yenmez, 2020) .
+However, these Pareto improvements have inherent limits: they cannot address DA’s poor rank distributions or student segregation (Ortega et al., 2025).
 Scope of Efficiency Gains.
@@
 (2024) show that DA is Pareto-inefficient with high probability in large i.i.d.
-random markets, and that the envy digraph typically contains a giant strongly connected component, implying that most students can potentially benefit from disjoint trading cycles.
+random markets and that most students can potentially benefit from disjoint trading cycles.
 Furthermore, all efficient mechanisms that dominate DA tend to improve nearly every student.
-This finding provides context for our analysis.
+These findings provide context for our analysis.
 While their paper shows that improvement mechanisms converge in their scope in large i.i.d.
-random markets, our paper demonstrates that this convergence does not hold in specific instances.
-We show that significant gaps can emerge between mechanisms in terms of both the number of students improved and the stability costs incurred.
+random markets, our paper demonstrates that this equivalence does not hold in specific instances, showing that the number of students who benefit can vary dramatically between different efficient mechanisms that dominate DA.
 3 Model
-Following Abdulkadiroğlu and Sönmez (2003), a school choice problem $P$ consists of a set of $n$ students $I$ and a set of schools $S$.
+A school choice problem $P$ consists of a set of $n$ students $I$ and a set of schools $S$.
 Each student $i\in I$ has a strict preference relation $\succ_{i}$ over the schools.
@@
 For a given school choice problem $P$, a matching $\mu$ is a mapping from $I$ to $S$ such that no school is matched to more students than its quota.
-We denote by $\mu_{i}$ the school to which student $i$ is assigned.
+We denote by $\mu_{i}$ the school to which student $i$ is assigned in $\mu$.
 The function $\textup{rk}_{i}:S\rightarrow\{1,\ldots,n\}$ specifies the rank of school $s$ according to the preference relation $\succ_{i}$ of student $i$:
@@
 so that student $i$’s most preferred school gets a rank of $1$. With some abuse of notation, we use the same rank function to specify the students’ rank per the priority profile of schools.
-Properties.
 A matching $\mu$ weakly Pareto-dominates matching $\nu$ if, for every student $i\in I$, $\textup{rk}_{i}(\mu_{i})\leq\textup{rk}_{i}(\nu_{i})$.
@@
 We say that student $j$ violates student $i$’s priority at school $s$ in matching $\mu$ if $i$ desires $s$, $\mu_{j}=s$, and $\textup{rk}_{s}(i)<\textup{rk}_{s}(j)$.
-If so, we call ($i,j,s$) and ($i,s$) a blocking triplet and a blocking pair at matching $\mu$, respectively.
-A matching $\mu$ is non-wasteful if every school $s$ that is desired by some student in $\mu$ satisfies $|\mu^{-1}_{s}|=q_{s}$.
+In this case, we say that student $i$’s envy towards $j$ is justified, and we call ($i,j,s$) and ($i,s$) a blocking triplet and a blocking pair in matching $\mu$, respectively.
+We use $B(\mu)$ to denote the set of blocking pairs in matching $\mu$.
+A matching $\mu$ is non-wasteful if every school $s$ that is desired by some student in $\mu$ satisfies $|\{i\in I:\mu_{i}=s\}|=q_{s}$.
 A matching $\mu$ is stable if it is non-wasteful and no student’s priority at any school is violated in $\mu$.
-Mechanisms.
-A mechanism is a function that maps every school choice problem to a matching. We particularly focus on three mechanisms.
-  1. 1.
-    Deferred Acceptance (DA).
-The student-proposing DA mechanism (Gale and Shapley, 1962) works as follows.
+Mechanisms with and without Consent.
+A mechanism is a function that maps every school choice problem to a matching.
+We focus on two well-known mechanisms.
+The first of these is the student-proposing Deferred Acceptance mechanism (DA, Gale and Shapley 1962), which works as follows.
 Students apply to their most preferred school.
@@
 The algorithm terminates when no new rejections occur, yielding the student-optimal stable matching $\mathtt{DA}(P)$.
-  2. 2.
-    Efficiency-Adjusted DA (EADA). To execute EADA, first we define an interrupter, which is a student-school pair $(i,s)$ such that:
+The second mechanism of interest is Top Trading Cycles on top of DA (DA+TTC).
+This mechanism first computes DA, then applies Gale’s Top Trading Cycles algorithm where each student initially “owns” their DA assignment (Shapley and Scarf, 1974).
+Students point to owners of their most preferred schools, and all students in cycles trade.
+The process repeats until no cycles remain.
+Standard mechanisms simply return a matching for any school choice problem.
+A consent-based mechanism $M$ instead maps every school choice problem and a subset of students $W\subseteq I$ to a matching.
+The interpretation of the subset of students $W$ is that these students agree to have their priorities violated at some schools as long as consenting to waive their priorities does not harm them, and thus we refer to $W$ as the consent structure.
+Formally, the property that waiving their priorities never harms consenting students means that for any problem $P$ and any consent structure $W\subseteq I\setminus\{i\}$:
+$\textup{rk}_{i}[M(P,W\cup\{i\})]\leq\textup{rk}_{i}[M(P,W)]$
+Note that for any fixed consent structure $W$, a consent-based mechanism $M(\cdot,W)$ reduces to a standard mechanism that maps problems to matchings.
+Efficiency-Adjusted Deferred Acceptance (EADA, Kesten 2010) is an example of a consent-based mechanism.
+To describe EADA, we first define an interrupter, which is a student-school pair $(i,s)$ such that:
   • student $i$ is tentatively accepted at $s$ in step $t$,
@@
   • student $i$ is rejected from school $s$ in step $t^{\prime\prime}\geq t^{\prime}$
-    To compute EADA, we follow the following procedure:
-    1. (a)
-      Run DA, find the last interrupter ($i,s$), in which student $i$ consents to waive his priority at school $s$.
-    2. (b)
-      Eliminate $s$ from the list of preferences of student $i$ and re-run DA.
-    3. (c)
-      Find the last consenting interrupter. If any, go to step 2. If there are no interrupters, the last DA allocation is implemented.
-    When all interrupter students consent to waive their priorities, EADA is Pareto-efficient.
-In that case, EADA can alternatively be computed by running DA, finding schools that did not reject anyone, matching them permanently to their DA match and removing such students and schools, and re-executing DA on the smaller problem until all schools are rejected (Tang and Yu, 2014).
-  3. 3.
-    Top Trading Cycles on top of DA (DA+TTC) This mechanism first computes DA, then applies Gale’s Top Trading Cycles algorithm where each student initially "owns" their DA assignment (Shapley and Scarf, 1974).
-Students point to owners of their most preferred schools, and all students in cycles trade.
-The process repeats until no cycles remain.
+We say that an interrupter $(i,s)$ consents if $i\in W$.
+Having defined the notion of an interrupter, we now describe the celebrated (and slightly more complex) EADA consent-based mechanism in detail below.
+Algorithm 1 Efficiency-Adjusted Deferred Acceptance (EADA)
+1:Input: Problem $P$ and consent structure $W\subseteq I$
+2:Compute initial matching $\mu^{0}=\mathtt{DA}(P)$
+3:while consenting interrupters exist do
+4:  Find the last round $t$ where a consenting interrupter $(i,s)$ was rejected
+5:  Identify all consenting interrupters $(i,s)$ in round $t$
+6:  if no consenting interrupters found then
+7:   break
+8:  end if
+9:  for each consenting interrupter $(i,s)$ in round $t$ do
+10:   Remove school $s$ from student $i$’s preferences
+11:  end for
+12:  Rerun DA with modified preferences
+13:end while
+14:return final matching
+We denote by $\mathtt{EADA}(P,W)$ the matching generated by Efficiency-Adjusted Deferred Acceptance algorithm in school choice problem $P$ with consent structure $W$.
+We refer to the mechanism $\mathtt{EADA}(\cdot,I)$ as EADA with full consent.
+EADA with full consent can be more straightforwardly computed by identifying every school who did not reject any student in DA’s execution (these schools are called under-demanded).
+After finding DA’s allocation, those schools are permanently matched with their assigned students, and such under-demanded schools and corresponding students are removed from DA.
+Then DA is executed in the simplified problem, and this procedure is repeated until all schools have been removed.
+This procedure was proposed by Tang and Yu (2014) and can be extended to partial consent structures, but this will not be necessary for our purposes.
+Setwise and Cardinally Minimally Unstable Mechanisms.
+Since it is well-known that every mechanism that Pareto-dominate DA generates blocking pairs in general, we need a way to compare their relative performance in terms of instability (e.g.
+blocking pairs).
+We use the notion of setwise minimally unstable mechanisms, a popular way of comparing the blocking pairs generated by assignment mechanisms (Tang and Zhang, 2021; Doğan and Ehlers, 2021a; Abdulkadiroğlu et al., 2020; Kwon and Shorrer, 2020; Bonkoungou and Nesterov, 2025).
+Definition 1.
+A Pareto-efficient mechanism $M$ is setwise minimally unstable if there does not exist any problem $P$ and any other Pareto-efficient mechanism $N$ such that $B(N(P))\subsetneq B(M(P))$.
+EADA with full consent satisfies this property (Doğan and Ehlers, 2021a; Tang and Zhang, 2021), as do many other mechanisms, since the setwise criterion creates only a partial order across mechanisms.
+Therefore, mechanisms that fail to be setwise minimally unstable are quite disappointing from a legal and fairness perspective, as they generate an unnecessarily large number of students who have their priorities violated and who may legally challenge their allocation.
+However, as we will demonstrate, this desirable stability property comes at a significant cost in terms of the number of students who can be improved.
+A different way to compare the instability among mechanisms that dominate DA is to count the number of blocking pairs they generate, rather than comparing them by set inclusion.
+This leads to a notion of cardinally more stable mechanisms.
+Definition 2.
+A Pareto-efficient mechanism $M$ is cardinally more stable than a Pareto-efficient mechanism $N$ if, for every problem $P$, $|B(M(P))|\leq|B(N(P))|$, with strict inequality for some $P$.
+We also define two other useful properties of mechanisms used throughout the paper.
+First, we say that a mechanism $M$ Pareto-dominates DA if: (i) for any problem $P$, the matching $M(P)$ weakly Pareto-dominates the matching $\mathtt{DA}(P)$, and (ii) there exists some problem $P$ where the matching $M(P)$ strictly Pareto-dominates the matching $\mathtt{DA}(P)$.
+Second, we say that a mechanism $M$ is Pareto-efficient if, for any problem $P$, the matching $M(P)$ is Pareto-efficient.
 Envy Digraphs and Improvement Cycles.
-To analyse improvements over DA, we use the envy digraph proposed by Ortega et al. (2024), which illustrates who envies whom after the initial DA allocation (envy need not be justified).
-Definition 1.
-The envy digraph $G^{\mathtt{DA}(P)}$ is a directed graph where:
-  • Nodes represent students in $I$.
-  • A directed edge $(i,j)$ exists if student $i$ envies student $j$’s assignment under DA, i.e. if $\textup{rk}_{i}(\mathtt{DA}_{j}(P))<\textup{rk}_{i}(\mathtt{DA}_{i}(P))$.
-We denote the edge $(i,j)$ as $i\rightarrow j$.
-A cycle in the digraph is a sequence of nodes $(i_{0}\to i_{1}\to\ldots\to i_{k}\to i_{0})$ such that there is a directed edge between each pair of consecutive nodes and no edge is repeated.
+To analyse improvements over DA, we adopt a graph-theoretic framework.
+Formally, the envy digraph $G^{\mathtt{DA}(P)}$ is a directed graph where nodes represent students and a directed edge $(i,j)$ exists if student $i$ envies student $j$’s assignment under DA.We denote the edge $(i,j)$ as $i\rightarrow j$.
+A cycle $C$ in the envy digraph is a sequence of nodes $C=(i_{0}\to i_{1}\to\ldots\to i_{k}\to i_{0})$ such that there is a directed edge between each pair of consecutive nodes and no edge is repeated.
 A trading cycle is a cycle in which every node appears exactly once, except for $i_{0}$.
 Unless specifically stated otherwise, every cycle mentioned henceforth is a trading cycle.
-As argued in Section 2, DA is rarely Pareto-efficient.
-Therefore, we focus on school choice problems where $G^{\mathtt{DA}(P)}$ contains at least one cycle.
-To systematically identify and implement improvements over DA, we use the concept of a feedback set.
-Definition 2.
+We focus on school choice problems where $G^{\mathtt{DA}(P)}$ contains at least one cycle.
 A feedback set $F$ is a collection of pairwise-disjoint cycles in $G^{\mathtt{DA}(P)}$ such that, when all nodes in these cycles are removed (together with their adjacent edges), the resulting subgraph is acyclic.
-[^2] Feedback node sets are sometimes called feedback vertex sets (Karp, 2010).
-The concept of a “cycle cover” is related, but requires that every node must be part of a cycle.
-The problem of choosing the maximum feedback set is sometimes known as cycle packing.
+[^3] Feedback sets, sometimes called feedback vertex sets, have been extensively studied in graph theory (Karp, 2010).
 Feedback sets represent sets of disjoint trading cycles that can be simultaneously implemented to improve upon DA.
@@
 If $i\in V(F)$, we say that $F$ covers $i$.
-A feedback set $F^{\star}$ is maximum if, for any other feedback set $F^{\prime}$, we have $|V(F^{\prime})|\leq|V(F^{\star})|$.
-Given a feedback set $F$, we define the matching $\mu^{F}$ as the allocation obtained when the trades in each cycle of $F$ are implemented, starting from DA. Formally:
-  • For student $i\in V(F)$ belonging to cycle $C$, $\mu^{F}_{i}=\mathtt{DA}_{\text{succ}(i)}(P)$, where $\text{succ}(i)$ denotes the successor of $i$ in cycle $C\in F$.
-  • For student $i\notin V(F)$, $\mu^{F}_{i}=\mathtt{DA}_{i}(P)$.
+Given a feedback set $F$, we define the matching $\mu^{F}$ as the allocation obtained when the trades in each cycle of $F$ are implemented, starting from DA.
 Our analysis builds upon a well-known result which connects improvements over DA to cycles in envy digraphs:
 Lemma 1 (Tang and Yu (2014)).
-Every Pareto improvement over DA corresponds to a set of trading cycles in $G^{\mathtt{DA}(P)}$.
-[^3] Similar arguments have been previously made in the literature by Abdulkadiroğlu et al.
-(2009); Kesten (2010), Erdil (2014) and Ortega et al.
-(2024).
-Having established the required framework, we are now ready to state our results.
+Every Pareto improvement over DA corresponds to a set of pairwise-disjoint trading cycles in $G^{\mathtt{DA}(P)}$.
+Having established our framework, we are now ready to state our results.
 4 Limitations of Existing Mechanisms
-How many students do existing mechanisms like EADA and DA+TTC actually improve compared to the maximum improvement possible over DA?
-To quantify this gap, we define the improvement ratio of a mechanism as follows.
-Let $\mathcal{I}(M,P)$ denote the set of students improved by the mechanism $M$ that Pareto-dominates DA in problem $P$:
-$\mathcal{I}(M,P)\coloneqq\{i\in I:\textup{rk}_{i}(M_{i}(P))<\textup{rk}_{i}(\mathtt{DA}_{i}(P))\}$ (2)
-We define the maximum improvement $\mathcal{I}(\mu^{*},P)$ as:
-$\mathcal{I}(\mu^{*},P)\coloneqq\max_{\mu\in\mathcal{M}^{\mathtt{DA}}(P)}|\mathcal{I}(\mu,P)|$ (3)
+How many students do existing mechanisms actually improve compared to the maximum improvement possible over DA?
+To quantify this gap, we define the improvement ratio of a mechanism.
+Let $\mathcal{I}(\mu,P)$ denote the set of students improved by the matching $\mu$ that Pareto-dominates DA in problem $P$:
+$\mathcal{I}(\mu,P)\coloneqq\{i\in I:\textup{rk}_{i}(\mu_{i})<\textup{rk}_{i}(\mathtt{DA}_{i}(P))\}$ (2)
+Let $\mathcal{M}^{\mathtt{DA}}(P)$ denote the set of matchings that weakly Pareto-dominate DA in $P$.
+We define the maximum improvement $\mathcal{I^{*}}(P)$ as:
+$\mathcal{I^{*}}(P)\coloneqq\max_{\mu\in\mathcal{M}^{\mathtt{DA}}(P)}|\mathcal{I}(\mu,P)|$ (3)
 Let $\mathcal{P}_{n}$ denote the class of all school choice problems with $n$ students. For any mechanism $M$ that Pareto-dominates DA, we define its improvement ratio as:
-$\rho(M;n)\coloneqq\max_{P\in\mathcal{P}_{n}}\frac{|\mathcal{I}(\mu^{*},P)|}{|\mathcal{I}(M,P)|}$ (4)
-For example, an improvement ratio of 2 means that the maximum mechanism may generate twice as many improvements as mechanism $M$.
-Our first result shows that this ratio can grow linearly with the size of the problem.
-[^4] Throughout Sections 4-4.1, we assume all students consent to waive their priorities, as even with full consent we obtain the worst possible bounds.
-We examine partial consent explicitly in Section 4.2 and formalize arbitrary consent structures in Section 5.
+$\rho(M;n)\coloneqq\max_{P\in\mathcal{P}_{n}}\frac{|\mathcal{I^{*}}(P)|}{|\mathcal{I}(M(P),P)|}$ (4)
+For example, an improvement ratio of 2 means that the mechanism that improves as many students as possible compared to DA may generate twice as many improvements as mechanism $M$.
+Our first result shows that this ratio can grow linearly with the size of the problem, and particularly that the improvement ratio of DA+TTC and EADA with full consent achieve the worst possible improvement ratio among all mechanisms that weakly Pareto-dominate DA.
 Theorem 1.
-For every $n>4$ there exists a problem $P$ such that the maximum improvement is $n-1$; while that of EADA and DA+TTC are both exactly 2:
-$\rho(\text{EADA};n)=\rho(\text{DA+TTC};n)=\frac{n-1}{2}$
+For any $n>4$ and any Pareto-efficient mechanism $M$ that Pareto-dominates DA:
+$\rho(\mathtt{EADA};n)=\rho(\mathtt{DA+TTC};n)=\frac{n-1}{2}\geq\rho(M;n)$
 Proof.
 Upper Bound.
-First we prove that the improvement ratio cannot exceed $\frac{n-1}{2}$.
-For the numerator: At most $n-1$ can improve over DA, as there is always at least one unimprovable student (Tang and Yu, 2014).
-For the denominator: The minimum number of students improved by EADA and DA+TTC is 2, because a trading cycle requires at least two students.
+First we prove that the improvement ratio cannot exceed $\frac{n-1}{2}$ for any mechanism that Pareto-dominates DA.
+For the numerator: At most n1n-1 students can improve over DA, as there is always at least one unimprovable student (Tang and Yu, 2014).
+For the denominator: The minimum number of students improved by any Pareto-efficient mechanism that Pareto-dominates DA is 2, because a trading cycle requires at least two students.
 [^4] We have assumed that DA is inefficient; otherwise, the minimum would be 0 and the ratio undefined.
 Therefore, the ratio cannot exceed $\frac{n-1}{2}$.
-Tightness. We construct a school choice problem with $n$ students and schools (all with unit capacity) where this improvement ratio is achieved for an arbitrary value of $n>4$.
+Tightness.
+To show this bound is achievable, we construct a school choice problem with $n$ students and schools (all with unit capacity) where EADA with full consent and DA+TTC achieve this improvement ratio while a different matching improves n1n-1 students.
 Let school priorities be such that, for every school $s_{k}$ except $s_{1}$:
@@
 The DA allocation appears in bold.
-The allocation that maximises the number of improvements appears in circles, whereas the matching recommended by both EADA and DA+TTC appears in squares (whenever different from DA).
+The allocation that maximises the number of improvements appears in circles, whereas the matching recommended by both EADA with full consent and DA+TTC appears in squares (whenever different from DA).
 Example 1
@@
 $s_{6}$ $s_{3}$ $s_{4}$ $s_{5}$ $s_{2}$ $s_{2}$ $\bm{s_{7}}$ $i_{7}$ $i_{1}$ $i_{2}$ $i_{3}$ $i_{4}$ $i_{5}$ $i_{6}$
-$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $\bm{s_{5}}$ $\bm{s_{6}}$cdot $i_{2}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
-$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
+$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $\bm{s_{5}}$ $\bm{s_{6}}$ $\cdot$ $i_{2}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
+$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
 We now demonstrate that the unique stable matching assigns each student $i_{k}$ to the school with the corresponding index $s_{k}$.
@@
 The uniqueness of the set of stable matchings can be verified by observing that the school-proposing DA algorithm yields the same outcome, with each school immediately being matched to its highest-priority student.
-Analysis of EADA.
-Under EADA, we begin by identifying under-demanded schools.
+Analysis of EADA with full consent.
+We begin by identifying under-demanded schools.
 First, $s_{n}$ is under-demanded and is removed together with $i_{n}$.
@@
 These students form a trading cycle where $i_{1}$ improves to $s_{2}$ and $i_{2}$ improves to $s_{1}$.
-Thus, EADA improves exactly 2 students (assuming $i_{n}$ consents to waive his priority at $s_{1}$, otherwise no student improves).
+Thus, EADA with full consent improves exactly 2 students (assuming $i_{n}$ consents to waive his priority at $s_{1}$, otherwise no student improves).
 Analysis of DA+TTC.
@@
 Maximum Improvement.
-There exists, however, a matching $\mu^{*}$ that dominates DA, obtained by implementing the cycles $(i_{1}\rightarrow i_{n-1}\rightarrow i_{1})$ and $(i_{2}\rightarrow i_{3}\rightarrow\ldots\rightarrow i_{n-2}\rightarrow i_{2})$, which improves $n-1$ students.
+There exists, however, a matching $\mu^{*}$ that dominates DA, obtained by implementing the cycles $(i_{1}\rightarrow i_{n-1}\rightarrow i_{1})$ and $(i_{2}\rightarrow i_{3}\rightarrow\ldots\rightarrow i_{n-2}\rightarrow i_{2})$, which improves n1n-1 students.
 Under this matching, student $i_{n}$ receives the same allocation as under DA, while all other students receive a more preferred school.
-Therefore, the improvement ratio equals $\frac{|I(\mu^{*},P)|}{|I(M,P)|}=\frac{n-1}{2}$ for any $n>4$, establishing the tightness of our bound. ∎
-Three remarks are in order now.
+Therefore, the improvement ratio equals $\frac{n-1}{2}$ for any $n>4$, establishing the tightness of our bound. ∎
+Some remarks are now in order.
 First, Kesten (2010) already pointed out in his seminal paper that larger improvements than EADA were possible.
-Our contribution here is to quantify how much larger those improvements can be, leading to an improvement gap that grows linearly with the problem size.
-Second, in our constructed example, the maximum improvement also dominates on rank positions improved: the 2-student cycle generates 4 total rank improvements while the $(n-1)$-student cycle generates $n$ total rank improvements.
-For $n\geq 6$, maximizing student count yields strictly higher rank improvement, showing that existing mechanisms can simultaneously underperform on both distributional and rank-efficiency dimensions, albeit this need not be the case in general.
-Third, in our example, the maximum improvement comes at the cost of generating more blocking pairs.
-In the particular example with seven agents, the maximum improvement generates at least three blocking pairs ($\{i_{1},s_{2}\},\{i_{2},s_{1}\},\{i_{7},s_{1}\}$) versus a single blocking pair generated by EADA and DA+TTC ($\{i_{7},s_{1}\}$).
-We do not take a stance on whether the larger improvement is worth the additional justified envy, our point is to make explicit that such reduction in justified envy comes at a high cost in terms of how many students are benefited.
-Furthermore, as we will show next, the larger improvement can sometimes lead to fewer blocking pairs.
+Our contribution here is to quantify how much larger those improvements can be even when all students consent to waive their priorities, leading to an improvement gap that grows linearly with the problem size.
+Second, in our example, the maximum improvement comes at the cost of generating more blocking pairs.
+In the particular example with seven agents, the larger improvement generates at least three blocking pairs ($\{i_{1},s_{2}\},\{i_{2},s_{1}\},\{i_{7},s_{1}\}$) versus a single blocking pair generated by EADA with full consent and DA+TTC ($\{i_{7},s_{1}\}$).
+We do not take a stance on whether the larger improvement is worth the additional justified envy.
+Our point is to make explicit that such reduction in justified envy comes at a high cost in terms of how many students are benefited.
+Furthermore, as we will show in the next subsection, larger improvements can sometimes lead to fewer blocking pairs.
+Third, it is natural to ask whether EADA with full consent and DA+TTC are the only mechanism that fails to implement larger improvements that benefit more students.
+We answer this question in the negative in the next Subsection.
+4.1 A General Impossibility Result
+We now generalize Theorem 1 to a wide class of mechanisms beyond EADA with full consent and DA+TTC, namely all those that are setwise minimally unstable, i.e.
+those which never produce a set of blocking pairs that is strictly larger by set inclusion than that generated by another Pareto-efficient mechanism that Pareto-dominates DA.
+Theorem 2.
+Among Pareto-efficient mechanisms that Pareto-dominate DA, every setwise minimally unstable mechanism achieves an improvement ratio of $\frac{n-1}{2}$.
+Proof.
+We use the same problem as in the proof of Theorem 1. In the case with seven students and schools, there are six cycles in DA’s envy digraph, summarized in Table 1 below.
+Table 1: Cycles and Blocking Pairs in Example 1
+Cycles Blocking Pairs
+$C_{1}=(i_{1}\to i_{2}\to i_{1})$ $(i_{7},s_{1})$
+$C_{2}=(i_{1}\to i_{2}\to i_{3}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
+$C_{3}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
+$C_{4}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
+$C_{5}=(i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2})$ $(i_{1},s_{2})$
+$C_{6}=(i_{1}\to i_{6}\to i_{1})$ $(i_{7},s_{1}),(i_{2},s_{1}),(i_{5},s_{6})$
+$C_{7}=(i_{1}\to i_{6}\to i_{2}\to i_{1})$ $(i_{7},s_{1}),(i_{1},s_{2}),(i_{5},s_{6})$
+Any Pareto-efficient mechanism that Pareto-dominates DA must implement at least one of these cycles (potentially two if mutually disjoint).
+In particular, the larger improvement that includes as many students as possible implements $C_{5}=(i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2})$ and $C_{6}=(i_{1}\to i_{6}\to i_{1})$.
+The cycles $C_{2},C_{3},C_{4},C_{6},C_{7}$ all generate a set of blocking pairs that is strictly larger (by set inclusion) than that generated by $C_{1}$.
+The cycle $C_{5}$ is incomparable to $C_{1}$, but implementing it by itself does not lead to a Pareto-efficient allocation, since the cycle $C_{6}$ could be jointly implemented.
+But if both $C_{5}$ and $C_{6}$ are implemented, the set of blocking pairs becomes $\{(i_{1},s_{2}),(i_{2},s_{1}),(i_{7},s_{1})\}$, a strict superset of the set of blocking pairs generated by $C_{1}$.
+Thus, we conclude that every setwise minimally unstable must implement $C_{1}$, therefore missing the larger improvements dictated by $C_{5}$ and $C_{6}$ jointly, leading to an improvement ratio of $\frac{n-1}{2}=\frac{6}{2}$.
+Finally, the extension to arbitrary $n$ is straightforward: there are $n$ cycles, with $C_{k}=(i_{1}\to\ldots\to i_{k+1}\to i_{1})$ for all integer $1\leq k\leq n-3$, all of them generating the blocking pair ($i_{n},s_{1}$), and only $C_{1}$ introducing exactly this blocking pair.
+There are exactly three more cycles: $C_{n-2}=(i_{2}\to i_{3}\to\ldots\to i_{n-3}\to i_{2})$, $C_{n-1}=(i_{1}\to i_{6}\to i_{1})$, and $C_{n}=(i_{1}\to i_{n-1}\to i_{2}\to i_{1})$.
+Only one collection of cycles that leaves the graph acyclic is setwise minimally unstable: $C_{1}$.
+However, the cycles $C_{n-2}$ and $C_{n-1}$ are disjoint (and thus compatible) and lead to an allocation where n1n-1 students improve, and therefore any setwise minimally unstable mechanism incurs in the largest improvement ratio possible, namely $\frac{n-1}{2}$.
+∎
+Noting that EADA is setwise minimally unstable (as proven by Tang and Zhang (2021) and Doğan and Ehlers (2021a)) makes it evident that Theorem 2 implies Theorem 1.
+Nonetheless, we present both Theorems independently to simplify the argument and improve their readability.
 4.2 Larger Improvements with Less Justified Envy
-Theorem 1 shows that EADA and DA+TTC can improve significantly fewer students, establishing a tight bound.
-In the example used in the proof, we showed that larger improvements came at a cost of more blocking pairs.
-Now we will show that this need not always be the case: in some problems, EADA and DA+TTC can be doubly dominated: they may improve fewer students while generating more justified envy.
-Proposition 1.
-There exist school choice problems where a larger improvement than EADA’s and DA+TTC’s is possible that generates fewer blocking pairs.
+Theorems 1 and 2 shows that EADA with full consent and DA+TTC can improve significantly fewer students compared to the DA baseline, but that such smaller coverage is unavoidable if we insist on mechanisms that are setwise minimally unstable.
+Now we show that the impossibility disappears if we compare blocking pairs simply by counting them, i.e.
+we consider cardinal minimal instability.
+To be precise, let $M$ and $N$ be Pareto-efficient mechanisms that Pareto-dominate DA.
+We say that $M$ improves more students than $N$ if for every problem $P$, we have $|\mathcal{I}(N(P),P)|\leq|\mathcal{I}(M(P),P)|$, with strict inequality for some $P$.
+We are now ready to state our result.
+Theorem 3.
+There exists a Pareto-efficient mechanism that Pareto-dominates DA that improves more students than EADA with full consent while being cardinally more stable than EADA with full consent.
+An analogous result holds for DA+TTC.
 Proof.
-Comparison with EADA.
+Comparison with EADA with full consent.
 Consider the following school choice problem with seven students and seven schools with one seat each.
@@
 $i_{1}$ $i_{2}$ $i_{3}$ $i_{4}$ $i_{5}$ $i_{6}$ $i_{7}$ $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $s_{6}$ $s_{7}$
-$s_{6}$ $s_{1}$ $s_{6}$ $s_{5}$ $s_{3}$ $s_{4}$ $s_{4}$ $\bm{i_{1}}$ $\bm{i_{2}}$ $\bm{i_{3}}$ $\bm{i_{4}}$ $\bm{i_{5}}$ $\bm{i_{6}}$cdot
-$s_{4}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $s_{6}$ $\boldsymbol{s_{6}}$ $\boldsymbol{s_{7}}$ $i_{5}$ $i_{1}$ $i_{5}$ $i_{7}$ $i_{4}$ $i_{3}$cdot
-$s_{2}$cdot $\cdot$ $\cdot$ $s_{4}$cdot $\cdot$ $i_{2}$cdot $i_{1}$ $i_{1}$ $i_{1}$ $i_{5}$cdot
-$s_{3}$cdot $\cdot$ $\cdot$ $s_{1}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$cdot $i_{1}$cdot
-$s_{5}$cdot $\cdot$ $\cdot$ $\bm{s_{5}}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{5}$cdot $\cdot$ $\cdot$
-$\bm{s_{1}}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
-The DA envy digraph appears below, showing the trading cycles implemented by EADA (Figure 1, improving 4 students through the cycle $i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$) and the maximum improvement (Figure 2, improving 6 students via two disjoint cycles, namely $i_{1}\to i_{2}\to i_{1}$ and $i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3}$).
-Interestingly, EADA generates 3 blocking pairs ($\{i_{3},s_{6}\},\{i_{5},s_{6}\},\{i_{7},s_{4}\}$), whereas the maximum improvement generates only two blocking pairs ($\{i_{1},s_{4}\},\{i_{7},s_{4}\}$).
+$s_{6}$ $s_{1}$ $s_{6}$ $s_{5}$ $s_{3}$ $s_{4}$ $s_{4}$ $\bm{i_{1}}$ $\bm{i_{2}}$ $\bm{i_{3}}$ $\bm{i_{4}}$ $\bm{i_{5}}$ $\bm{i_{6}}$ $\cdot$
+$s_{4}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $s_{6}$ $\boldsymbol{s_{6}}$ $\boldsymbol{s_{7}}$ $i_{5}$ $i_{1}$ $i_{5}$ $i_{7}$ $i_{4}$ $i_{3}$ $\cdot$
+$s_{2}$ $\cdot$ $\cdot$ $\cdot$ $s_{4}$ $\cdot$ $\cdot$ $i_{2}$ $\cdot$ $i_{1}$ $i_{1}$ $i_{1}$ $i_{5}$ $\cdot$
+$s_{3}$ $\cdot$ $\cdot$ $\cdot$ $s_{1}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$ $\cdot$ $i_{1}$ $\cdot$
+$s_{5}$ $\cdot$ $\cdot$ $\cdot$ $\bm{s_{5}}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{5}$ $\cdot$ $\cdot$ $\cdot$
+$\bm{s_{1}}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
+The DA envy digraph appears below, showing the trading cycles implemented by EADA with full consent (Figure 1), improving 4 students through the cycle $(i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1})$ and an alternative, larger improvement (Figure 2), benefiting 6 students via two disjoint cycles, namely $(i_{1}\to i_{2}\to i_{1})$ and $(i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3})$.
+Interestingly, EADA with full consent generates 3 blocking pairs ($\{i_{3},s_{6}\},\{i_{5},s_{6}\},\{i_{7},s_{4}\}$), whereas the alternative, larger improvement generates only two blocking pairs ($\{i_{1},s_{4}\},\{i_{7},s_{4}\}$).
 [^5] Note that here we are simply counting the number of blocking pairs; there is a literature (see Section 2) arguing that that some are more relevant than others.
-Note that the maximum improvement improves every improvable student, because $i_{7}$ is unimprovable since nobody envies him.
+Note that the larger improvement benefits every student who could benefit under some allocation that Pareto-dominates DA, because $i_{7}$ cannot trade with anyone since no student envies him.
 $i_{1}$$i_{2}$$i_{3}$$i_{4}$$i_{5}$$i_{6}$$i_{7}$$i_{1}$$i_{4}$$i_{5}$$i_{6}$
-Figure 1: EADA
+Figure 1: EADA, full consent.
 $i_{1}$$i_{2}$$i_{3}$$i_{4}$$i_{5}$$i_{6}$$i_{7}$$i_{1}$$i_{2}$$i_{3}$$i_{4}$$i_{5}$$i_{6}$
-Figure 2: Maximum Improvement
+Figure 2: Larger improvement.
+To conclude the proof, consider the mechanism $M$ that chooses the same matching as EADA with full consent in every problem except the aforementioned one.
+Such matching clearly is Pareto-efficient, Pareto-dominates DA, improves more students than EADA with full consent and is cardinally more stable than EADA.
 Comparison with DA+TTC.
-To show that a larger improvement can generate fewer blocking pairs than DA+TTC and can be implemented with fewer students consenting, consider the following example with five students and five schools with unit capacity.
+To show that a larger improvement can generate fewer blocking pairs than DA+TTC, consider the following example with five students and five schools with unit capacity.
 DA’s allocation appears in bold.
@@
 $s_{2}$ $s_{1}$ $s_{1}$ $s_{1}$ $s_{1}$ $\bm{i_{1}}$ $\bm{i_{2}}$ $\bm{i_{3}}$ $\bm{i_{4}}$ $\bm{i_{5}}$
-$s_{4}$ $s_{3}$ $s_{2}$ $s_{2}$ $\bm{s_{5}}$ $i_{5}$ $i_{4}$ $i_{4}$ $i_{5}$cdot
-$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$cdot $i_{4}$ $i_{3}$ $i_{2}$ $i_{1}$cdot
-$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{2}$ $i_{1}$cdot $\cdot$ $\cdot$
+$s_{4}$ $s_{3}$ $s_{2}$ $s_{2}$ $\bm{s_{5}}$ $i_{5}$ $i_{4}$ $i_{4}$ $i_{5}$ $\cdot$
+$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $\cdot$ $i_{4}$ $i_{3}$ $i_{2}$ $i_{1}$ $\cdot$
+$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{2}$ $i_{1}$ $\cdot$ $\cdot$ $\cdot$
 In DA+TTC, students $i_{1}$ and $i_{2}$ trade their DA allocations, generating four blocking pairs: $(i_{4},s_{1}),(i_{5},s_{1}),(i_{3},s_{2}),(i_{4},s_{2})$.
-In contrast, the maximum improvement is the cycle $(i_{1}\to i_{4}\to i_{1}),(i_{2}\to i_{3}\to i_{2})$, generating only one blocking pair: $(i_{5},s_{1})$.
+In contrast, the largest improvement is the cycle $(i_{1}\to i_{4}\to i_{1}),(i_{2}\to i_{3}\to i_{2})$, generating only one blocking pair: $(i_{5},s_{1})$.
+To conclude the proof, consider the mechanism $M$ that chooses the same matching as DA+TTC in every problem except the aforementioned one.
+Such matching clearly is Pareto-efficient, Pareto-dominates DA, improves more students than DA+TTC and is cardinally more stable than DA+TTC.
 ∎
-While DA+TTC, and in particular, EADA, choose improvement cycles taking priorities into account, therefore satisfying several weak stability properties (see Section 2), Proposition 1 shows that they may generate doubly dominated matchings in some problems—improving fewer students while creating more blocking pairs—since alternative matchings can achieve better outcomes on both dimensions simultaneously.
-Noting that the larger improvements in the previous two examples are obtained by DA+TTC and EADA, respectively, yields the following incomparability corollary.
+We now discuss some implications of Theorem 3.
+First, our two examples reveal an interesting asymmetry: in Example 2, DA+TTC achieves the larger improvement that doubly dominates EADA, whereas in Example 3, EADA with full consent produces the improvement that double dominates DA+TTC.
+This observation yields the following incomparability result:
 Corollary 1.
-EADA and DA+TTC are incomparable in terms of size of improvement and number of blocking pairs, i.e. there exist problems where DA+TTC doubly dominates EADA and vice versa.
-We find Corollary 1 interesting because DA+TTC has been considered suboptimal regarding a number of axioms (Kesten, 2010; Troyan et al., 2020), yet it is not uniformly doubly dominated by the more well-regarded EADA mechanism.
-4.2 Consent and Improvement Size
-In his seminal paper, Kesten (2010) argues that “the likelihood of getting stuck without gaining much Pareto improvement clearly increases with the cycle size".
-This intuition seems natural—coordinating more students should be harder.
-However, this relationship is more nuanced than originally thought.
-Let us recall that, in EADA, a student agrees to waive his priorities before the DA mechanism is executed at schools where he is an interrupter.
-Kesten shows that such students never benefit nor gain from giving their consent.
-Let us divide students who consent into two classes: non-beneficiary consenters are those who neither benefit nor are harmed by giving their consent, and beneficiary consenters who explicitly benefit from the trading cycles generated by their consent to waive their priorities.
-It is reasonable to conjecture that beneficiary consents are more likely to be obtained than non-beneficiary ones, as the former ones directly benefit rather than providing consent altruistically.
-With this distinction in mind, let us revisit Example 2 in the proof of Proposition 1.
-EADA implements the four-cycle ($i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$), requiring non-beneficiary consent from both $i_{3}$ (at $s_{6}$) and $i_{7}$ (at $s_{4}$), as well as beneficiary consent from $i_{5}$ (at $s_{6}$).
-In contrast, the maximum improvement implements two disjoint cycles requiring fewer non-beneficiary consents (1 versus 2 of EADA), and the same number of beneficiary consents (one each, see Table 1).
-If non-beneficiary consents, which rely purely on altruism, fail, EADA cannot improve any student, and yet the largest improvement can still execute the two-cycle ($i_{1},i_{2},i_{1}$), given the consent of potential beneficiary $i_{5}$.
-Student $i_{5}$ does not benefit given the lack of non-beneficiary consents, but since consents are expressed ex-ante this does not affect his decision on whether to consent or not.
-Table 1: Consent Requirements in Example 2
-Mechanism Consent Required Improvement Size
-Non-beneficiary Beneficiary
-EADA $i_{3},i_{7}$ $i_{5}$ 4 ($i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$)
-$i_{4}$, $i_{3}$ (no $i_{7}$) 2 ($i_{1}\to i_{5}\to i_{1}$)
-$i_{5}$ (no $i_{7},i_{4}$) 2 ($i_{1}\to i_{2}\to i_{1}$)
-otherwise 0
-Maximum $i_{7}$ $i_{1}$ 6 ($i_{1}\to i_{2}\to i_{1}$), ($i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3}$)
-Improvement $i_{5}$ 2 ($i_{1}\to i_{2}\to i_{1}$)
-otherwise 0
-We summarize this observation as Proposition 2 below.
-Proposition 2.
-There exist problems for which a Pareto improvement that strictly improves more students than EADA requires fewer total and non-beneficiary consents.
-Furthermore, in some problems, larger improvements can be implemented relying solely on beneficiary consents, whereas EADA may fail to implement smaller improvements without non-beneficiary consent.
-Having examined the potential shortcomings of existing mechanisms, we proceed to analyse a new mechanism that improves as many students as possible compared to the DA baseline.
-5 Maximum Implementable Improvement over Deferred Acceptance
-Our theoretical analysis has revealed that existing mechanisms can fall short of maximizing the number of students who benefit from efficiency enhancements to DA.
-The Maximum Implementable Improvement over Deferred Acceptance (MIIDA) mechanism addresses this limitation by identifying and implementing improvements that benefit the maximum possible number of students given the consent structure.
-To define MIIDA formally, let us incorporate consent into the mathematical framework.
-Each student reports strict preferences over schools $\succ_{i}$ and whether they consent $c_{i}=1$ (or not, with $c_{i}=0$) to waive their priority at schools.
-A consent-based mechanism maps a school choice problem and a consent decision by each student to a matching.
-Let $C\subseteq I$ denote the set of students who agree to waive their priority when they do not benefit from the proposed trades.
-For any edge $i\to j$ in the envy digraph $G^{\mathtt{DA}(P)}$, let $\text{consent}(i\to j)$ denote the set of students whose consent is required for student $i$ to receive student $j$’s DA assignment.
-Specifically,
-$\text{consent}(i\to j)=\{k\in I:k\text{ has higher priority than }i\text{ at school }\mathtt{DA}_{j}(P)\}$
-For any feedback set $F$, we define $L(F)=\bigcup_{C\in F}\bigcup_{(i\to j)\in C}\text{consent}(i\to j)$ as the set of all students whose consent is required to implement the trades in $F$.
-A feedback set $F$ is implementable given consent set $C$ if $L(F)\setminus V(F)\subseteq C$, so that all required consents come from either beneficiaries of the trades or students in the consent set $C$.
-We define MIIDA as follows.
-Algorithm 1 Maximum Improvement over Deferred Acceptance (MIIDA)
-1:Compute the student-optimal stable matching $\mu^{0}=\mathtt{DA}(P)$
-2:Set iteration counter $t=0$
-3:repeat
-4:     Generate the envy digraph $G^{\mu^{t}}$ for current matching $\mu^{t}$
-5:     if $G^{\mu^{t}}$ contains no cycles then
-6:         break
-7:     end if
-8:     Let $N^{t}=I\setminus C$ be the set of non-consenting students
-9:     Remove nodes $N^{t}$ and adjacent edges from $G^{\mu^{t}}$ to obtain reduced graph $\tilde{G}^{\mu^{t}}$
-10:     if $\tilde{G}^{\mu^{t}}$ contains no cycles then
-11:         break $\triangleright$ No improvements possible with available consent
-12:     end if
-13:     Find maximum feedback set $F^{t}$ in $\tilde{G}^{\mu^{t}}$ (with a fixed tie-breaking rule)
-14:     Implement trades in $F^{t}$ to produce matching $\mu^{t+1}$
-15:     $t\leftarrow t+1$
-16:until convergence
-17:return final matching $\mu^{t}$
-MIIDA iteratively finds maximum implementable feedback sets in the envy digraph until no further improvements are possible.
-Iteration is essential because implementing trades changes the envy structure: after students exchange assignments, new envy relationships emerge while others disappear.
-To illustrate, consider the consent-labeled envy digraph from our earlier example.
-MIIDA proceeds as follows:
-  1. 1.
-    In the first iteration, MIIDA identifies the maximum feedback set
-    $F={(i_{1}\to i_{2}\to i_{1}),(i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3})}$
-    covering 6 students, which requires consent from $L(F)\setminus V(F)={i_{7}}$.
-  2. 2.
-    If $\{i_{1},i_{7}\}\subset C$, MIIDA implements $F$.
-  3. 3.
-    If the required consents are not available, MIIDA seeks the largest implementable feedback set in the graph. If $i_{7}\in C$, it finds $F^{\prime}={(i_{1}\to i_{2}\to i_{1})}$ covering 2 students.
-  4. 4.
-    If $i_{5}\in C$, implement $F^{\prime}$. Otherwise, make no changes over DA.
-MIIDA satisfies a number of important desiderata.
-Theorem 2.
-MIIDA has the following properties:
-  1. 1.
-    Weakly Pareto-dominates DA.
-  2. 2.
-    Is Pareto-efficient if all students consent.
-  3. 3.
-    Only violates the priorities of students who consent.
-  4. 4.
-    Maximises the number of students who benefit from Pareto improvements over DA given the consent structure (so it is never doubly-dominated).
-  5. 5.
-    Can generate strictly fewer, but also strictly more, blocking pairs than EADA and DA+TTC, depending on the school choice problem.
-  6. 6.
-    No student has incentives not to consent.
-Specifically, for every student consenting is a very weakly dominant strategy in the complete information preference revelation game induced by MIIDA in which students simultaneously report their preferences and whether they consent to waive their priorities.
-[^7] A strategy is very weakly dominant for a player if, for every strategy profile, it gives a weakly higher utility than every other strategy (e.g.
-Chen and Micali, 2015).
-On the other hand, a weakly dominant strategy additionally requires that for some strategy profile it must yield a strictly higher utility than any other strategy available to the player.
-  7. 7.
-    Can be computed in polynomial time.
-Proof.
-  1. 1.
-    Pareto-dominance: Follows directly from MIIDA’s construction: it implements only cycles where each participant receives a strictly preferred school relative to DA (by envy digraph definition).
-Non-participants keep their DA assignments, guaranteeing weak Pareto dominance for everyone.
-  2. 2.
-    Pareto-efficient if all consent: When $C=I$, $\tilde{G}^{\mu^{t}}=G^{\mu^{t}}$. Termination occurs only when no cycles exist, implying no further Pareto improvements are possible (by Lemma 1)).
-  3. 3.
-    Priority violations require consent: For any edge $i\to j$ implemented, consent is required from all $k$ with higher priority at $s=\mathtt{DA}_{j}(P)$.
-The condition $L(F^{t})\setminus V(F^{t})\subseteq C$ ensures only consenting non-beneficiaries’ priorities are violated.
-  4. 4.
-    Maximizes improvements given consent: Let $S$ be students improved by MIIDA.
-Suppose another mechanism $M$ improves $T$ with $|T|>|S|$.
-By Lemma 1, $T$ corresponds to disjoint cycles $\mathcal{C}_{T}$ in $G^{\mathtt{DA}(P)}$ implementable under $C$.
-In MIIDA’s first iteration, it selects a feedback set $F^{0}$ maximizing $|V(F^{0})|$ over implementable sets.
-Since $\mathcal{C}_{T}$ is implementable: $|V(F^{0})|\geq|T|$.
-But $S$ includes $V(F^{0})$ (at minimum), so $|S|\geq|T|$, contradiction.
-Thus no such $M$ exists.
-  5. 5.
-    Variable blocking pairs: Proven by Theorem 1 and Proposition 1: Example 1 shows MIIDA can generate more blocking pairs than EADA/DA+TTC; Example 2 shows it can also generate fewer.
-  6. 6.
-    Consent incentives: Fix student $i\in I$, any preference profile $P=(\succ^{\prime}_{i},\succ^{\prime}_{-i})$, and any consent decisions $c_{-i}$ by other students.
-Let $\mu^{\mathtt{DA}(P)}=\mathtt{DA}(P)$ and let $C=\{j\neq i:c_{j}=1\}\cup\{i\}$ denote the consent set when student $i$ consents.
-Let $F$ denote the maximum implementable feedback set under consent set $C$.
-    When $c_{i}=0$, by MIIDA’s construction, $i$ cannot participate in any trading cycle. Therefore $\mathtt{MIIDA}_{i}(P;c_{i}=0,c_{-i})=\mu^{\mathtt{DA}(P)}_{i}$.
-    When $c_{i}=1$:
-    • If $i\notin V(F)=\mathtt{MIIDA}_{i}(P;c_{i}=1,c_{-i})=\mu^{\mathtt{DA}(P)}_{i}$
-    • If $i\in V(F)=\mathtt{MIIDA}_{i}(P;c_{i}=1,c_{-i})=\mu^{\mathtt{DA}(P)}_{\text{succ}(i)}$
-    where $\text{succ}(i)$ is $i$’s successor in the cycle containing $i$ within $F$.
-If $i\in V(F)$, then by definition of envy digraph $G^{\mu^{\mathtt{DA}(P)}}$, there exists edge $i\to\text{succ}(i)$, which means $\mu^{\mathtt{DA}(P)}_{\text{succ}(i)}\succ^{\prime}_{i}\mu^{\mathtt{DA}(P)}_{i}$.
-Therefore $\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{\text{succ}(i)})<\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{i})$.
-    It follows that
-    $\textup{rk}_{i}(\mathtt{MIIDA}(P;c_{i}=1,c_{-i}))\leq\textup{rk}_{i}(\mu^{\mathtt{DA}(P)}_{i})=\textup{rk}_{i}(\mathtt{MIIDA}(P;c_{i}=0,c_{-i}))$
-    with strict inequality when $i\in V(F)$. Since this holds for arbitrary $(\succ^{\prime}_{i},\succ^{\prime}_{-i},c_{-i})$, consenting is a very weakly dominant strategy for student $i$.
-  7. 7.
-    Polynomial runtime: MIIDA runs in polynomial time through its iterative structure.
-DA is computed once initially in $O(n^{2})$ time (Gale and Shapley, 1962).
-Each subsequent iteration involves: (i) generating the envy digraph for the current matching in $O(n^{2})$ time, and (ii) finding a maximum feedback set.
-The cycle packing problem can be solved in $O(m\sqrt{n})$ time where $m$ is the number of edges and $n$ is the number of vertices (Biro et al., 2009).
-Since the envy digraph has at most $n^{2}$ edges, this step takes $O(n^{2}\sqrt{n})=O(n^{2.5})$ time per iteration.
-Since each iteration covers at least one additional student in trading cycles, and at most $n-1$ students can be improved over DA (Tang and Yu, 2014), MIIDA terminates within $n$ iterations.
-Therefore, the overall complexity is $O(n^{3.5})$.
-While MIIDA improves as many students as possible given the consent structure available, and is never doubly dominated, it also has a number of important limitations.
-Arguably, the most severe is its lack of stability.
-Since it chooses cycles without taking into account priorities, it can lead to more blocking pairs than EADA and DA+TTC (but also fewer, as we have seen).
-In fact, MIIDA sometimes generates a set of blocking pairs that is a strict superset of EADA’s blocking pairs.
-Unfortunately, this is unavoidable by any mechanism that does not miss a significant number of improvement cycles.
-To state this result formally, let us denote by $B(M,P)$ the set of blocking pairs generated by mechanism $M$ in problem $P$.
-We say mechanism $M$ is setwise justified envy minimal if for any problem $P$, there is no alternative mechanism $M^{\prime}$ such that $B(M^{\prime},P)\subsetneq B(M,P)$.
-A number of mechanisms, including EADA (Doğan and Ehlers, 2021a), satisfy this partial comparison property.
-However, setwise justified envy minimality is at odds with larger improvements.
-Theorem 3.
-Among Pareto-efficient mechanisms that weakly Pareto-dominate DA, every mechanism that is setwise justified envy minimal achieves an improvement ratio of $\frac{n-1}{2}$.
-Proof.
-Recall Example 1 in the proof of Theorem 1. In the case with $7$ students and schools, there are six cycles in DA’s envy digraph, summarized in Table 2.
-Table 2: Cycles and Blocking Pairs in Example 1
-Cycles Blocking Pairs
-$K_{1}=(i_{1}\to i_{2}\to i_{1})$ $(i_{7},s_{1})$
-$K_{2}=(i_{1}\to i_{2}\to i_{3}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
-$K_{3}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
-$K_{4}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
-$K_{5}=(i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2})$ $(i_{1},s_{2})$
-$K_{6}=(i_{1}\to i_{6}\to i_{1})$ $(i_{7},s_{1}),(i_{2},s_{1}),(i_{5},s_{6})$
-$K_{7}=(i_{1}\to i_{6}\to i_{2}\to i_{1})$ $(i_{7},s_{1}),(i_{1},s_{2}),(i_{5},s_{6})$
-Any efficient mechanism that weakly dominates DA must implement at least one of these cycles (potentially two if mutually disjoint).
-In particular, MIIDA implements ($i_{1}\to i_{6}\to i_{1}$) and ($i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2}$).
-The cycles $K_{2},K_{3},K_{4},K_{6},K_{7}$ all generate a set of blocking pairs that is strictly larger (by set inclusion) than that generated by $K_{1}$.
-The cycle $K_{5}$ is incomparable to $K_{1}$, but implementing it by itself does not lead to a Pareto-efficient allocation, since the cycle $K_{6}$ could be jointly implemented.
-But if both $K_{5}$ and $K_{6}$ are implemented, the set of blocking pairs becomes $\{(i_{1},s_{2}),(i_{2},s_{1}),(i_{7},s_{1})\}$, a strict superset of the set of blocking pairs generated by $K_{1}$.
-Thus, we conclude that every setwise justified envy minimal mechanism must implement $K_{1}$therefore missing the larger improvements dictated by $K_{5}$ and $K_{6}$ jointly, leading to an improvement ratio of $\frac{n-1}{2}=\frac{6}{2}$.
-Finally, the extension to arbitrary $n$ is straightforward: there are $n$ cycles, with $K_{k}=(i_{1}\to\ldots\to i_{k+1}\to i_{1})$ for all integer $31\leq k\leq n-$ 3, all of them generating the blocking pair ($i_{n},s_{1}$), and only $K_{1}$ introducing exactly this blocking pair.
-There are exactly three more cycles: $K_{n-2}=(i_{2}\to i_{3}\to\ldots\to i_{n-3}\to i_{2})$, $K_{n-1}=(i_{1}\to i_{6}\to i_{1})$, and $K_{n}=(i_{1}\to i_{n-1}\to i_{2}\to i_{1})$.
-Only one collection of cycles that leaves the graph acyclic is setwise justifiable envy minimal: $K_{1}$.
-However, the cycles $K_{n-2}$ and $K_{n-1}$ are compatible and lead to an allocation where $n-1$ students improve, and therefore any setwise justified envy minimal mechanism incurs in the largest improvement ratio possible, namely $\frac{n-1}{2}$.
-∎
-MIIDA also has important limitations beyond its lack of stability properties.
-The mechanism is manipulable and may have multiple solutions.
-Nonetheless, despite its limitations, we believe MIIDA serves as a useful mechanism to help us better understand the tradeoffs between efficiency and stability when improving over the celebrated DA mechanism.
+EADA with full consent and DA+TTC are incomparable in terms of size of improvement and number of blocking pairs, i.e. there exist problems where DA+TTC doubly dominates EADA and vice versa.
+We find Corollary 1 useful because DA+TTC has been considered suboptimal regarding several axioms (Kesten, 2010; Troyan et al., 2020), yet it is not consistently doubly dominated by the more well-regarded EADA mechanism.
+Second, the possibility of simultaneously achieving broader student coverage and fewer blocking pairs suggests that alternative efficient mechanisms that Pareto-dominate DA merit investigation.
+In particular, understanding the properties of such mechanisms—their computational complexity, laboratory performance, and incentive properties—represents an exciting area for future research.
+While EADA has deservedly received wide praise in the literature, our results point to unexplored opportunities within the class of mechanisms that Pareto-dominate DA that could consistently outperform EADA in this particular criterion of more beneficiaries with fewer blocking pairs.
+Finally, we discuss the relationship between the impossibility result (Theorem 2) and the possibility result (Theorem 3).
+Both setwise and cardinal minimal instability capture legitimate concerns about mechanism design, though they reflect different policy goals.
+Setwise minimal instability focuses on the diversity of potential challenges to the allocation, recognizing that each blocking pair represents a distinct grievance that could motivate legal or political opposition.
+Cardinal minimal instability, by contrast, adopts a more utilitarian perspective that treats each blocking pair as imposing similar costs regardless of the specific students involved.
+While our impossibility result applies specifically to setwise minimally unstable mechanisms, this does not invalidate the setwise criterion but rather illuminates the fundamental trade-offs it entails.
+How much this limitation matters in practice depends critically on policymaker objectives: those prioritizing broad political consensus and legal robustness may find the setwise criterion compelling despite its constraints on student coverage, while those focused on maximizing total improvements or minimizing aggregate unfairness may prefer cardinal comparisons.
 5 Conclusion
-We have shown that EADA and DA+TTC can be doubly dominated, improving as few as 2 students when $n-1$ could benefit, while generating more blocking pairs than larger improvements.
-MIIDA addresses this by maximizing the number of students who benefit, though our impossibility theorem proves this necessarily comes at the cost of losing EADA’s minimal justified envy guarantee.
-An equally promising alternative would modify how EADA handles consent to minimize blocking pairs rather than maximize improvements.
-While such a mechanism might not improve as many students as MIIDA, it could nonetheless avoid double-dominance.
-This represents a complementary approach to navigating the efficiency-stability frontier that warrants future investigation.
+Our analysis establishes that setwise minimally unstable mechanisms, including EADA, face fundamental constraints on student coverage.
+We prove these mechanisms achieve the worst possible improvement ratio, potentially helping only $2$ students when n1n-1 could benefit.
+However, this limitation disappears under cardinal comparisons of blocking pairs: alternative mechanisms can simultaneously improve more students than EADA while generating fewer blocking pairs.
+Our findings reveal that the choice between setwise and cardinal minimal instability represents an important design decision.
+While setwise minimally unstable mechanisms offer protection against diverse legal challenges, it does so at the cost of significantly reduced student participation in efficiency gains.
+The relatively underexplored class of mechanisms that Pareto-dominate DA contains promising alternatives for policymakers willing to prioritize broader student coverage over setwise stability constraints.